### v1

# Positive definiteness and Fell bundles over discrete groups

Abstract
We introduce a natural concept of positive definiteness for bundle maps between Fell bundles over (possibly different) discrete groups and describe several examples. Such maps induce completely positive maps between the associated full cross-sectional C*-algebras in a functorial way. Under the assumption that the kernel of the homomorphism connecting the groups under consideration is amenable, they also induce completely positive maps between the associated reduced cross-sectional C*-algebras. As an application, we define an approximation property for a Fell bundle over a discrete group which generalizes Exel’s approximation property and still implies the weak containment property; both approximation properties coincide when the unit fibre is nuclear.
MSC 2020: 46L55 (Primary); 43A35, 46L07, 46M18 (Secondary)
Keywords: Fell bundle, cross-sectional $C^{*}$-algebra, positive definite bundle map, completely positive map, approximation property, nuclearity
1 Introduction
The concept of positive definiteness for complex-valued functions on a group $G$ has a long history (it goes back to Toeplitz in 1911 when $G={\mathbb{Z}}$ and to Mathias in 1923 when $G={\mathbb{R}}$; for more details, see [59] and references therein). Its importance is ubiquitous in various areas of mathematics, in particular in connection with the development in the 1940’s and 50’s of the representation theory of locally compact groups and of the associated group $C^{*}$-algebras (see e.g. [22, 28, 29]). On the other hand, completely positive linear maps play a prominent rôle in the modern theory of operator algebras and in quantum information theory (cf. e.g. [56, 14, 33]). Haagerup discovered in 1978 that there is a very useful link between the two concepts. Namely, he proved in [31, Theorem 3.1] that given an action $\alpha$ of a locally compact group $G$ on a von Neumann algebra ${\cal M}$, every continuous positive definite function $\psi$ on $G$ gives rise to a normal completely positive linear map $M_{\psi}$ on the crossed product ${\cal M}\rtimes_{\alpha}G$ such that $M_{\psi}$ acts on each $g$-component by multiplication with $\psi(g)$. This link also holds in the context of group $C^{*}$-algebras and group von Neumann algebras, as noted by Haagerup in his seminal paper [32] and exploited in later works, such as [21] where continuous positive definite functions on $G$ are identified as multipliers of the Fourier algebra $A(G)$ of $G$. In the setting of discrete unital $C^{*}$-dynamical systems, a similar link was established in a series of paper [8, 9, 10], where broader concepts of multipliers were introduced. As a sample of related articles we mention [50, 51, 52, 34, 6, 39, 49].
Given two $C^{*}$-algebras $A$ and $B$, it is often the case that one needs to construct completely positive linear maps from $A$ to $B$. When $A$ and $B$ belong to some class of $C^{*}$-algebras, one may hope to find some specific procedure achieving this. In this paper, we consider this problem for the class $\mathcal{F}$ (resp. $\mathcal{F}_{\rm r}$) of $C^{*}$-algebras which may be described as the full (resp. reduced) cross-sectional $C^{*}$-algebra $C^{*}({\mathcal{A}})$ (resp. $C_{r}^{*}({\mathcal{A}})$) of a Fell bundle ${\mathcal{A}}$ over a discrete group, as studied in [26, 27] (and e.g. in [45, 40]; see also [28, 29] for a more general treatment dealing with $C^{*}$-algebraic bundles, i.e., Fell bundles over locally compact groups). The class $\mathcal{F}$ (resp. $\mathcal{F}_{\rm r}$) is large as it contains all full (resp. reduced) twisted group $C^{*}$-algebras, and more generally all full (resp. reduced) $C^{*}$-crossed products associated to discrete twisted $C^{*}$-dynamical systems, even those for which the action of the group is only assumed to be partial [25]. Also, every $C^{*}$-algebra on which there exists a maximal (resp. normal) nondegenerate coaction by a discrete group belongs to $\mathcal{F}$ (resp. $\mathcal{F}_{\rm r}$), see [23, Proposition 4.2] (resp. [58, Corollary 3.9]). In particular, both classes include any $C^{*}$-algebra on which there exists a nondegenerate coaction by a discrete amenable group, e.g., any $C^{*}$-algebra on which there exists an action of a compact abelian group, such as Cuntz algebras and, more generally, (higher-rank) graph $C^{*}$-algebras. Note here that one may always choose to regard a $C^{*}$-algebra $B$ as the cross-sectional $C^{*}$-algebra of a Fell bundle over the trivial group, hence that our set up includes the case where no assumption is put on $B$.
After a preliminary section (Section 2), our main findings are gathered in Section 3. To describe these, let us first assume that $A=C^{*}({\mathcal{A}})$ and $B=C^{*}({\mathcal{B}})$, where ${\mathcal{A}}$ is a Fell bundle over a discrete group $G$ and ${\mathcal{B}}$ is a Fell bundle over a discrete group $H$, and let $\varphi:G\to H$ be a homomorphism. Consider an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T:{\mathcal{A}}\to{\mathcal{B}}$, meaning that $T$ maps the fibre of ${\mathcal{A}}$ over any $g\in G$ into the fibre of ${\mathcal{B}}$ over $\varphi(g)$ in a linear and bounded way. Inspired by the definition of positive definiteness for a complex function on a group and the definition of complete positivity for a linear map between $C^{*}$-algebras, we introduce a natural notion of positive definiteness for such a bundle map, and show that $T$ is positive definite if and only if there exists a completely positive linear map $\Phi_{T}:A\to B$ which “extends” $T$ (see Theorem 3.10 for a more precise statement). Our proof relies on a reformulation of positive definiteness in terms of the matrix $C^{*}$ algebras associated to a Fell bundle by Abadie and Ferraro in [1], and builds on the Stinespring dilation theorem for Fell bundles due to Buss, Ferraro and Sehnem in [18, Appendix]. After introducing a category of Fell bundles over discrete groups whose arrows are pairs $(\varphi,T)$, where $\varphi\in{\rm Hom}(G,H)$ and $T$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map, we get that the association ${\mathcal{A}}\mapsto C^{*}({\mathcal{A}}),\,(\varphi,T)\mapsto\Phi_{T}$ is a functor into the category of $C^{*}$-algebras with completely positive linear maps as arrows.
Next, we consider the case where $A=C_{r}^{*}({\mathcal{A}})$ and $B=C_{r}^{*}({\mathcal{B}})$. It is not difficult to see that an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T$ has to be positive definite whenever there exists a completely positive linear map $M_{T}:A\to B$ which “extends” $T$. However, the converse statement can not hold in general. This may be deduced from a result in the recent book of Bekka and de la Harpe, cf. [11, Corollary 8.C.15], which says that the homomorphism $\varphi:G\to H$ extends to a *-homomorphism from $C_{r}^{*}(G)$ into $C_{r}^{*}(H)$ if and only if the kernel of $\varphi$ is an amenable subgroup of $G$. Using this result in combination with Fell’s absorption principle for Fell bundles and a generalization of another result of Buss, Ferraro and Sehnem ([18, Proposition 4.8]), we are able to show that the converse does hold if we also assume that $\ker(\varphi)$ is amenable, cf. Theorem 3.12. The association ${\mathcal{A}}\mapsto C_{r}^{*}({\mathcal{A}}),(\varphi,T)\mapsto M_{T}$ is then also functorial if one considers the subcategory of Fell bundles whose arrows satisfy that $\ker(\varphi)$ is amenable.
In Section 4 we illustrate the usefulness of Theorem 3.10 and Theorem 3.12 in a long series of examples, showing that these two results are not only generalizations of previous results of the same flavor, but also lead to new applications. In particular, we devote a section (Section 5) to discuss a new approximation property for a Fell bundle ${\mathcal{A}}$ over a discrete group. We call it the PD-approximation property since it requires the existence of a net of uniformly bounded, finitely supported positive definite bundle maps converging to the identity map on each fibre, in analogy with the well-known characterization of the amenability of a discrete group in terms of positive definite functions. The PD-approximation property for ${\mathcal{A}}$ is (formally) weaker than Exel’s approximation property [26, 27], but still implies the weak containment property (i.e., amenability of ${\mathcal{A}}$ in the sense of Exel), cf. Theorem 5.4. Moreover, when it is satisfied, then $C^{*}({\mathcal{A}})$ (resp. $C^{*}_{r}({\mathcal{A}})$) is nuclear if and only if the unit fibre of ${\mathcal{A}}$ is nuclear, cf. Theorem 5.6. Using current knowledge from [3], it follows that the PD-approximation property for ${\mathcal{A}}$ is equivalent to Exel’s approximation property when the unit fibre of ${\mathcal{A}}$ is assumed to be nuclear. However, the arguments used to deduce this fact give no clue whether it also holds when the unit fibre of ${\mathcal{A}}$ is not nuclear.
In the final section (Section 6) we discuss some possible further developments. Let us mention here another one. Abadie and Ferraro have introduced in [1] (see also [2]) the notion of a right Hilbert ${\mathcal{B}}$-bundle for a Fell bundle ${\mathcal{B}}$ over a group, which reduces to the notion a right Hilbert $C^{*}$-module when the group is trivial. In a forthcoming article we will consider left actions of Fell bundles on such right Hilbert bundles and discuss the connection with positive definite bundle maps. This will provide a useful tool to produce positive definite bundle maps. Moreover, we will also study how $C^{*}$-correspondences over the cross-sectional $C^{*}$-algebras arise from such actions.
When we were about to finish the first draft of the present article, an interesting preprint by Buss, Kwasniewski, McKee and Skalski appeared [19], where the existing theories of Fourier-Stieltjes algebras associated to groups [36], to groupoids [60, 54] and to discrete twisted unital $C^{*}$-dynamical systems [10] is extended to the category of twisted actions by étale groupoids on $C^{*}$-bundles. Only a few of their results deal with the general case of Fell bundles over étale groupoids, hence with general Fell bundles over discrete groups, and there is little overlap between our papers.
2 Preliminaries
Throughout this article, $G$ and $H$ will always denote discrete groups. The unit of any group will be denoted by $e$. We will be considering Fell bundles over discrete groups. Our main reference on this topic will be Exel’s book [27], and we will follow his notation and terminology. For the ease of the reader we give below a short introduction. We assume familiarity with Hilbert $C^{*}$-modules, and follow the notation in Lance’s book [46]. In particular, if $X$ is a right Hilbert module over a $C^{*}$-algebra $A$, ${\mathcal{L}}_{A}(X)$ will denote the $C^{*}$-algebra of adjointable operators on $X$. The left regular representation of $G$ on $\ell^{2}(G)$ will be denoted by $\lambda^{G}$.
Let ${\mathcal{A}}=(A_{g})_{g\in G}$ be a Fell bundle over $G$. So each fibre $A_{g}$ is a Banach space and we often use ${\mathcal{A}}$ to also denote the disjoint union of the $A_{g}$’s. There is a multiplication map ${\mathcal{A}}\times{\mathcal{A}}\to{\mathcal{A}},(a,b)\mapsto ab$, and an involutive map ${\mathcal{A}}\to{\mathcal{A}},a\mapsto a^{*}$, such that $A_{g}A_{h}\subseteq A_{gh}$ and $A_{g}^{*}=A_{g^{-1}}$ for all $g,h\in G$, which satisfy some natural conditions and turn the unit fibre $A_{e}$ into a $C^{*}$-algebra. The vector space $C_{c}({\mathcal{A}})$ consisting of all finitely supported functions $f:G\to{\mathcal{A}}$ such that $f(g)\in A_{g}\text{ for all }g\in G$, becomes a *-algebra with respect to the operations defined by
$(f_{1}\ast f_{2})(h)=\sum_{g\in G}f_{1}(g)f_{2}(g^{-1}h),\quad f^{*}(h)=f(h^{-1})^{*}$
for all $h\in G$. The space $C_{c}({\mathcal{A}})$ is also a right pre-Hilbert $A_{e}$-module, the right action of $A_{e}$ being defined pointwise and the inner product being given by
$\langle f_{1},f_{2}\rangle_{A_{e}}=\sum_{h\in G}f_{1}(h)^{*}f_{2}(h).$
By completion we obtain a right Hilbert $A_{e}$-module, denoted by $\ell^{2}({\mathcal{A}})$. The (left) regular representation $\lambda^{\mathcal{A}}=(\lambda^{\mathcal{A}}_{g})_{g\in G}$ of ${\mathcal{A}}$ in ${\mathcal{L}}_{A_{e}}(\ell^{2}({\mathcal{A}}))$ is determined for each $g\in G$ by
$(\lambda^{\mathcal{A}}_{g}(a)f)(h)=a\,f(g^{-1}h)\quad\text{whenever }a\in A_{g},f\in C_{c}({\mathcal{A}})\text{ and }h\in G.$
It induces a faithful *-representation $\iota^{\mathcal{A}}:C_{c}({\mathcal{A}})\to{\mathcal{L}}_{A_{e}}(\ell^{2}({\mathcal{A}}))$ given by
$\iota^{\mathcal{A}}(f)=\sum_{g\in G}\lambda^{\mathcal{A}}_{g}(f(g))\quad\text{ for all }f\in C_{c}({\mathcal{A}}).$
The reduced cross-sectional $C^{*}$-algebra $C_{r}^{*}({\mathcal{A}})$ is the $C^{*}$-subalgebra of ${\mathcal{L}}_{A_{e}}(\ell^{2}({\mathcal{A}}))$ generated by $\iota^{\mathcal{A}}(C_{c}({\mathcal{A}}))$, i.e., by $\{\lambda_{g}^{\mathcal{A}}(a):g\in G,a\in A_{g}\}$. The full cross-sectional $C^{*}$-algebra $C^{*}({\mathcal{A}})$ is defined as the $C^{*}$-completion of $C_{c}({\mathcal{A}})$ with respect to the universal $C^{*}$-norm $\|\cdot\|_{\rm u}$ given by
$\|f\|_{\rm u}=\sup\big{\{}p(f):p\text{ is a }C^{*}\text{-seminorm on }C_{c}({\mathcal{A}})\big{\}}.$
For $f\in C_{c}({\mathcal{A}})$, we set $\|f\|_{r}:=\|\iota^{\mathcal{A}}(f)\|$. We then have $\|f\|_{r}\leq\|f\|_{\rm u}$.
For $g\in G$, we let $j^{\mathcal{A}}_{g}:A_{g}\to C_{c}({\mathcal{A}})$ denote the canonical injection. If $a\in A_{g}$ we will sometimes write $a\odot g$ instead of $j^{\mathcal{A}}_{g}(a)$. Further, we let $\kappa^{\mathcal{A}}:C_{c}({\mathcal{A}})\to C^{*}({\mathcal{A}})$ denote the canonical injection and set $\widehat{j}^{\mathcal{A}}_{g}:=\kappa^{\mathcal{A}}\circ j^{\mathcal{A}}_{g}:A_{g}\to C^{*}({\mathcal{A}})$. For any $f\in C_{c}({\mathcal{A}})$, we have $f=\sum_{g\in G}j^{{\mathcal{A}}}_{g}(f(g))$, so
$\kappa^{\mathcal{A}}(f)=\sum_{g\in G}\widehat{j}^{\mathcal{A}}_{g}(f(g)).$
As each $\widehat{j}^{\mathcal{A}}_{g}$ is isometric, this gives that
$\|\kappa^{\mathcal{A}}(f)\|_{\rm u}\leq\sum_{g\in G}\|\widehat{j}^{\mathcal{A}}_{g}(f(g))\|_{\rm u}=\sum_{g\in G}\|f(g)\|.$ (1)
The canonical *-homomorphism $\Lambda^{\mathcal{A}}$ from $C^{*}({\mathcal{A}})$ onto $C_{r}^{*}({\mathcal{A}})$ is determined by
$\Lambda^{\mathcal{A}}\big{(}\,\widehat{j}_{g}^{\mathcal{A}}(a)\big{)}=\lambda_{g}^{\mathcal{A}}(a)$
for all $g\in G$ and $a\in A_{g}$, i.e., it satisfies $\iota^{\mathcal{A}}=\Lambda^{\mathcal{A}}\circ\kappa^{\mathcal{A}}$. The following diagram allows to visualize the various maps:
[…]
In addition, for each $g\in G$, there is a contractive linear map $E_{g}^{\mathcal{A}}:C_{r}^{*}({\mathcal{A}})\to A_{g}$ satisfying that for each $h\in G$ and $a\in A_{h}$, we have
$E_{g}^{\mathcal{A}}\big{(}\lambda^{\mathcal{A}}_{h}(a)\big{)}=\begin{cases}a,\quad\text{ if }g=h,\\ 0,\quad\text{ if}~g\neq h.\end{cases}$
For $x\in C_{r}^{*}({\mathcal{A}})$, $E_{g}^{\mathcal{A}}(x)$ may be thought of as the Fourier coefficient of $x$ at $g$. The map $E^{\mathcal{A}}:=E_{e}^{\mathcal{A}}$ is a faithful conditional expectation from $C_{r}^{*}({\mathcal{A}})$ onto $A_{e}$ (when one identifies $A_{e}$ with $\lambda_{e}^{\mathcal{A}}(A_{e})$).
We will need the following lemma several times. It follows readily from [27, Lemma 17.2]. For completeness, we give the proof.
Lemma 2.1.
Assume $a\in(A_{e})^{+}$ and $c\in{\mathcal{A}}$. Then $c^{*}ac\in(A_{e})^{+}$ and $c^{*}ac\leq\|a\|\,c^{*}c$. Hence,
$\|c^{*}ac\|~\,\leq\,\|a\|\,\|c\|^{2}.$
Proof.
Write $a=b^{*}b$ for some $b\in A_{e}$. Then $c^{*}ac=c^{*}b^{*}bc=(bc)^{*}bc\in(A_{e})^{+}$ and [27, Lemma 17.2] gives that
$c^{*}ac=~c^{*}b^{*}bc~\,\leq\,\|b\|^{2}\,c^{*}c=\|a\|\,c^{*}c.$
Hence $\|c^{*}ac\|\,\leq\,\|a\|\,\|c^{*}c\|=\|a\|\,\|c\|^{2}$. $\square$
We will also make use of the matrix C-algebras associated by Abadie and Ferraro to a Fell bundle ${\mathcal{A}}$ in [1]. We recall here their definition. Let $g_{1},\ldots,g_{n}\in G$ and set $\mathbf{g}:=(g_{1},\ldots,g_{n})\in G^{n}$. Then
$M_{\mathbf{g}}({\mathcal{A}}):=\big{\{}R=[r_{ij}]\in M_{n}({\mathcal{A}}):r_{ij}\in A_{g_{i}^{-1}g_{j}}\text{ for all }i,j=1,\ldots,n\big{\}}$
is a $C^{*}$-algebra with respect to the natural operations, the norm being inherited from the natural embedding of $M_{\mathbf{g}}({\mathcal{A}})$ into $M_{n}(C^{*}({\mathcal{A}}))$ (or $M_{n}(C_{r}^{*}({\mathcal{A}}))$). We note that $M_{\mathbf{g}}({\mathcal{A}})$ can also be represented by adjointable operators on a right Hilbert $A_{e}$-module. Indeed, consider the right Hilbert $A_{e}$-module $A_{\mathbf{g}}:=A_{g_{1}^{-1}}\oplus\cdots\oplus A_{g_{n}^{-1}}$ obtained by taking the direct sum of the $A_{g_{i}}$’s (considered as right Hilbert $A_{e}$-modules), whose inner product is given by
$\big{\langle}(a_{1},\ldots,a_{n}),(a^{\prime}_{1},\ldots,a^{\prime}_{n})\big{\rangle}_{A_{e}}=\sum_{i=1}^{n}a_{i}^{*}a^{\prime}_{i}$
for $a_{i},a^{\prime}_{i}\in A_{g_{i}^{-1}}$, $i=1,\ldots,n$. We can let each $R=[r_{ij}]\in M_{\mathbf{g}}({\mathcal{A}})$ act on $A_{\mathbf{g}}$ by
$L_{R}(a_{1},\ldots,a_{n})=(a^{\prime}_{1},\ldots,a^{\prime}_{n}),\ \text{ where }a^{\prime}_{i}:=\sum_{j=1}^{n}r_{ij}a_{j}\text{ for each }i=1,\ldots,n.$
It is easy to check that the operator $L_{R}:A_{\mathbf{g}}\to A_{\mathbf{g}}$ is adjointable with $(L_{R})^{*}=L_{R^{*}}$, and that the map $R\mapsto L_{R}$ from $M_{\mathbf{g}}({\mathcal{A}})$ into $\mathcal{L}_{A_{e}}(A_{\mathbf{g}})$ is a faithful *-homomorphism. Thus, the norm on $M_{\mathbf{g}}({\mathcal{A}})$ satisfies that $\|R\|=\|L_{R}\|$.
3 Positive definite bundle maps and complete positivity
In this section we let ${\mathcal{A}}=(A_{g})_{g\in G}$ be a Fell bundle over a group $G$, ${\mathcal{B}}=(B_{h})_{h\in H}$ be a Fell bundle over a group $H$ and $\varphi:G\to H$ be a group homomorphism, i.e., $\varphi\in{\rm Hom}(G,H)$.
3.1 Positive definite bundle maps
Definition 3.1.
An ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map is a family $T=(T_{g})_{g\in G}$ of maps $T_{g}:A_{g}\to B_{\varphi(g)}$ which are linear and bounded for all $g\in G$. Alternatively, we can think of $T$ as a map $T:{\mathcal{A}}\to{\mathcal{B}}$ such that for each $g\in G$ the restriction $T_{g}$ of $T$ to $A_{g}$ is a bounded linear map from $A_{g}$ into $B_{\varphi(g)}$.
Note that $T^{*}=(T^{*}_{g})_{g\in G}$, defined by
$T^{*}_{g}(a)=T_{g^{-1}}(a^{*})^{*}\text{~for all $g\in G$ and $a\in A_{g}$},$
is then also an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map.
An ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T$ is said to be self-adjoint if it satisfies that $T^{*}=T$, i.e.,
$T_{g}(a)^{*}=T_{g^{-1}}(a^{*})$
for all $g\in G$ and $a\in A_{g}$. Moreover, it is said to be multiplicative if it satisfies that
$T_{gg^{\prime}}(aa^{\prime})=T_{g}(a)T_{g^{\prime}}(a^{\prime})$
for all $g,g^{\prime}\in G$ and $a\in A_{g},a^{\prime}\in A_{g^{\prime}}$.
A pair $(\varphi,T)$ where $\varphi\in{\rm Hom}(G,H)$ and $T$ is a self-adjoint, multiplicative ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map will be called a (Fell bundle) morphism from ${\mathcal{A}}$ to ${\mathcal{B}}$.
In the case where $G=H$ and $\varphi={\rm id}_{G}$, so ${\mathcal{A}}$ and ${\mathcal{B}}$ are both Fell bundles over $G$, we will use the term ${\mathcal{A}}$$-{\mathcal{B}}$ bundle map instead of ${\mathcal{A}}$-${\rm id}_{G}$$-{\mathcal{B}}$ bundle map. Moreover, if ${\mathcal{A}}={\mathcal{B}}$, we will just say ${\mathcal{B}}$-bundle map instead of ${\mathcal{B}}$$-{\mathcal{B}}$ bundle map.
Remark 3.2.
Considering the case where $G=H$ and $\varphi={\rm id}_{G}$, Exel defines a morphism from ${\mathcal{A}}$ to ${\mathcal{B}}$ to be an ${\mathcal{A}}$$-{\mathcal{B}}$ bundle map $T$ which is self-adjoint and multiplicative, and he proves that it gives rise to a *-homomorphism from $C^{*}({\mathcal{A}})$ into $C^{*}({\mathcal{B}})$ (resp. from $C_{r}^{*}({\mathcal{A}})$ into $C_{r}^{*}({\mathcal{B}})$), cf. [27, Section 21.1]. We will later see (cf. Example 4.4) that in general, every morphism $(\varphi,T):{\mathcal{A}}\to{\mathcal{B}}$ gives rise to a *-homomorphism from $C^{*}({\mathcal{A}})$ into $C^{*}({\mathcal{B}})$; moreover, if we assume that ${\rm ker}(\varphi)$ is amenable, then we will also get a *-homomorphism from $C_{r}^{*}({\mathcal{A}})$ into $C_{r}^{*}({\mathcal{B}})$.
Definition 3.3.
We will say that an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ is positive definite whenever we have
$\sum_{i,j=1}^{n}b_{i}\,T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\,b_{j}^{*}\,\in(B_{e})^{+}$ (2)
for all $n\in{\mathbb{N}}$, $g_{1},\ldots,g_{n}\in G$ and all $a_{i}\in A_{g_{i}}$, $b_{i}\in B_{\varphi(g_{i})}$, $i=1,\ldots,n$.
Remark 3.4.
Let $T=(T_{g})_{g\in G}$ be an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. We note that $T$ is positive definite if and only if for every $n\in{\mathbb{N}}$, $\mathbf{g}=(g_{1},\ldots,g_{n})\in G^{n}$ and $a_{i}\in A_{g_{i}}$, $i=1,\ldots,n$, the matrix
$\Big{[}T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\Big{]}$
is positive in the $C^{*}$-algebra $M_{\varphi(\mathbf{g})}({\mathcal{B}})$, where $\varphi(\mathbf{g}):=(\varphi(g_{1}),\ldots,\varphi(g_{n}))\in H^{n}$.
Indeed, consider $\mathbf{g}=(g_{1},\ldots,g_{n})\in G^{n}$ and $a_{i}\in A_{g_{i}}$ for $i=1,\ldots,n$. Let then $S:=[s_{ij}]$ be the matrix in $M_{\varphi(\mathbf{g})}({\mathcal{B}})$ given by
$s_{ij}~:=T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\in B_{\varphi(g_{i}^{-1}g_{j})}=B_{\varphi(g_{i})^{-1}\varphi(g_{j})}$
for each $i,j\in\{1,\dots,n\}$. Then for $b_{1}\in B_{\varphi(g_{i})},\ldots,b_{n}\in B_{\varphi(g_{n})}$, we have that
$\sum_{i,j=1}^{n}b_{i}\,T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\,b_{j}^{*}=\sum_{i=1}^{n}b_{i}\,\Big{(}\sum_{j=1}^{n}s_{ij}\,b_{j}^{*}\Big{)}=\big{\langle}(b_{1}^{*},\ldots,b_{n}^{*}),L_{S}(b_{1}^{*},\ldots,b_{n}^{*})\big{\rangle}$
Using [46, Lemma 4.1], it follows from this equality that if $T$ is ${\mathcal{B}}$-positive definite, then $L_{S}$ is positive in $\mathcal{L}_{B_{e}}(B_{\varphi(\mathbf{g})})$. As the map $R\mapsto L_{R}$ is a faithful *-homomorphism from $M_{\varphi(\mathbf{g})}({\mathcal{B}})$ into $\mathcal{L}_{B_{e}}(B_{\varphi(\mathbf{g})})$, this is equivalent to $S$ being positive in $M_{\varphi(\mathbf{g})}({\mathcal{B}})$. This shows the forward implication. The converse implication can be shown by reversing the arguments above.
Proposition 3.5.
Let $T=(T_{g})_{g\in G}$ be a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. Then the following properties hold.
  • The map $T_{e}:A_{e}\to B_{e}$ is completely positive.
  • $T$ is self-adjoint.
  • For $g\in G$ and $a\in A_{g}$, we have  $T_{g}(a)^{*}T_{g}(a)\leq\|T_{e}\|\,T_{e}(a^{*}a)$((((in $B_{e}$)))).
Proof.
The first assertion follows readily by choosing all $g_{i}$’s to be equal to $e$ in the definition of positive definiteness. Next, pick an approximate unit $\{u_{i}\}_{i\in I}$ for $A$. Let $i\in I,g\in G,a\in A_{g}$. Choosing $g_{1}=e,g_{2}=g,a_{1}=u_{i}$ and $a_{2}=a$, and setting $\mathbf{g}=(e,g)$, we get that the matrix
$\begin{bmatrix}T_{e}(u_{i}^{2})&T_{g}(u_{i}a)\\ T_{g^{-1}}(a^{*}u_{i})&T_{e}(a^{*}a)\end{bmatrix}\in M_{\varphi(\mathbf{g})}({\mathcal{B}})^{+}.$
In particular, this matrix is self-adjoint, which gives that $T_{g}(u_{i}a)^{*}=T_{g^{-1}}(a^{*}u_{i})$.
Taking now the limit w.r.t. $i$ and using that $T_{g}$ is bounded, we get that $T_{g}(a)^{*}=T_{g^{-1}}(a^{*})$, and the second assertion follows.
Moreover, since $T_{e}(u_{i}^{2})$ and $T_{e}(a^{*}a)$ are positive elements in $A$, it follows from [46, Lemma 5.2 (iii)] that
$T_{g^{-1}}(a^{*}u_{i})T_{g}(u_{i}a)=T_{g}(u_{i}a)^{*}T_{g}(u_{i}a)\leq\|T_{e}(u_{i}^{2})\|\,T_{e}(a^{*}a)\leq\|T_{e}(u_{i})\|\,T_{e}(a^{*}a).$
Taking the limit w.r.t. $i$ and using again that $T_{g}$ is bounded, we get that the third assertion holds (since $\|T_{e}\|=\lim_{i}\|T_{e}(u_{i})\|$ as $T_{e}$ is completely positive). $\square$
Note that if ${\mathcal{A}}$ is unital (i.e., if $A_{e}$ is unital), replacing each $u_{i}$ with $1_{A_{e}}$, we don’t need to take any limit in the proof above, hence the boundedness assumption for each $T_{g}$ is not required in this case for the second and the third properties above to hold; the boundedness of each $T_{g}$ follows then readily from the third property, and could therefore be omitted in the definition of a positive definite bundle map.
As an immediate consequence of Proposition 3.5, we get
Corollary 3.6.
Let $T=(T_{g})_{g\in G}$ be a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. Then
$\sup_{g\in G}\|T_{g}\|=\|T_{e}\|.$
3.2 Completely positive maps
Our interest in positive definite bundle maps is due to their connection with completely positive maps. Here is the first evidence of this fact.
Proposition 3.7.
Let $M:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ be a completely positive linear map. For each $g\in G$, define a map $T_{g}:A_{g}\to B_{\varphi(g)}$ by
$T_{g}(a)=E^{\mathcal{B}}_{\varphi(g)}\big{(}M(\lambda^{\mathcal{A}}_{g}(a))\big{)}\quad\text{for each }a\in A_{g}.$
Then $T=(T_{g})_{g\in G}$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map.
Proof.
It is clear that each $T_{g}$ is linear and bounded, so $T$ is an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map.
Let $g_{1},\ldots,g_{n}\in G$, $a_{i}\in A_{g_{i}}$ and $b_{i}\in B_{\varphi(g_{i})}$ for $i=1,\ldots,n$. Then for each $i,j$ we have
$b_{i}\,T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\,b_{j}^{*}=b_{i}\,E^{\mathcal{B}}_{\varphi(g_{i}^{-1}g_{j})}\big{(}M(\lambda^{\mathcal{A}}_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\big{)}\,b_{j}^{*}=b_{i}\,E^{\mathcal{B}}_{\varphi(g_{i})^{-1}\varphi(g_{j})}\big{(}w_{ij}\big{)}\,b_{j}^{*}$
where $w_{i,j}:=M(\lambda^{\mathcal{A}}_{g_{i}}(a_{i})^{*}\lambda^{\mathcal{A}}_{g_{j}}(a_{j}))\in C_{r}^{*}({\mathcal{B}})$.
We note that the matrix $W:=[w_{ij}]\in M_{n}(C_{r}^{*}({\mathcal{B}}))$ is positive as $M$ is assumed to be completely positive. Considering the direct sum $(\ell^{2}({\mathcal{B}}))^{n}$ as a right Hilbert $B_{e}$-module, we let $M_{n}(C_{r}^{*}({\mathcal{B}}))$ act on $(\ell^{2}({\mathcal{B}}))^{n}$ in the natural way.
Using [27, Proposition 17.12] and setting $\Omega:=\big{(}j^{\mathcal{B}}_{\varphi(g_{1})}(b_{1})^{*},\ldots,j^{\mathcal{B}}_{\varphi(g_{n})}(b_{n})^{*}\big{)}\in(\ell^{2}({\mathcal{B}}))^{n}$, we therefore get that
$\sum_{i,j=1}^{n}b_{i}\,T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\,b_{j}^{*}=\sum_{i,j=1}^{n}\big{\langle}j^{\mathcal{B}}_{\varphi(g_{i})}(b_{i})^{*},w_{ij}\,j^{\mathcal{B}}_{\varphi(g_{j})}(b_{j})^{*}\big{\rangle}_{\ell^{2}({\mathcal{B}})}=\big{\langle}\Omega,W\,\Omega\big{\rangle}_{(\ell^{2}({\mathcal{B}}))^{n}}$
is positive in $B_{e}$. Thus, $T$ is positive definite. $\square$
Proceeding in a similar way, it should be straightforward for the reader to check that the following analogous result holds.
Proposition 3.8.
Let $\Phi:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ be a completely positive linear map. Then we get a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ by setting
$T_{g}(a)=E^{\mathcal{B}}_{\varphi(g)}\big{(}\Lambda^{\mathcal{B}}(\Phi(\,\widehat{j}^{\,{\mathcal{A}}}_{g}(a)))\big{)}\quad\text{for each }g\in G\text{~and }a\in A_{g}.$
Let $T=(T_{g})_{g\in G}$ be an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. We can then define a linear map $\phi_{T}:C_{c}({\mathcal{A}})\to C_{c}({\mathcal{B}})$ by setting
$\phi_{T}\big{(}f)=\sum_{g\in G}j^{\mathcal{B}}_{\varphi(g)}(T_{g}(f(g)))\quad\text{for all }f\in C_{c}({\mathcal{A}}).$
(Note that the sum above is finite since $f$ has finite support.) It is the only linear map from $C_{c}({\mathcal{A}})$ to $C_{c}({\mathcal{B}})$ satisfying that
$\phi_{T}\big{(}j_{g}^{\mathcal{A}}(a)\big{)}=j^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\quad\text{for all }g\in G\text{ and }a\in A_{g}.$
Definition 3.9.
Let $T=(T_{g})_{g\in G}$ be an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. We will say that $T$ is reduced ((((resp. full )))) if $\phi_{T}$ is bounded w.r.t. the reduced (resp. universal) $C^{*}$-norms on $C_{c}({\mathcal{A}})$ and $C_{c}({\mathcal{B}})$, and will denote the natural extension of $\phi_{T}$ to a linear bounded map from $C_{r}^{*}({\mathcal{A}})$ into $C_{r}^{*}({\mathcal{B}})$ (resp. from $C^{*}({\mathcal{A}})$ into $C^{*}({\mathcal{B}})$) by $M_{T}$ (resp. $\Phi_{T}$).
Thus, an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ is reduced when there exists a (uniquely determined) linear bounded map $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ satisfying that $M_{T}\circ\iota^{\mathcal{A}}=\iota^{\mathcal{B}}\circ\phi_{T}$, that is,
$M_{T}\Big{(}\sum_{g\in G}\lambda^{\mathcal{A}}_{g}(f(g))\Big{)}=\sum_{g\in G}\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(f(g)))\quad\text{ for all }~f\in C_{c}({\mathcal{A}}),$
[…]
Similarly, an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ is full when there exists a (uniquely determined) linear bounded map $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ satisfying that $\Phi_{T}\circ\kappa^{\mathcal{A}}=\kappa^{\mathcal{B}}\circ\phi_{T}$, that is,
$\Phi_{T}\Big{(}\sum_{g\in G}\,\widehat{j}^{\mathcal{A}}_{g}(f(g))\Big{)}=\sum_{g\in G}\,\widehat{j}^{\mathcal{B}}_{\varphi(g)}(T_{g}(f(g)))\quad\text{ for all }~f\in C_{c}({\mathcal{A}}),$
[…]
We note that if $T=(T_{g})_{g\in G}$ is a reduced ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map such that $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ is completely positive, then $T$ is positive definite. Indeed, this follows from Proposition 3.7 (with $M=M_{T}$), since
$T_{g}(a)=E^{\mathcal{B}}_{\varphi(g)}\big{(}\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\big{)}=E^{\mathcal{B}}_{\varphi(g)}\big{(}M_{T}(\lambda^{\mathcal{A}}_{g}(a))\big{)}$
for every $g\in G$ and $a\in A_{g}$.
We also note that if $T$ is a full ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map such that $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ is completely positive, then $T$ is positive definite. Indeed, for every $g\in G$ and $a\in A_{g}$, we have
$\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))=\Lambda^{\mathcal{B}}(\,\widehat{j}^{\,{\mathcal{B}}}_{\varphi(g)}(T_{g}(a)))=\Lambda^{\mathcal{B}}(\Phi_{T}(\,\widehat{j}^{\,{\mathcal{A}}}_{g}(a))),$
thus
$T_{g}(a)=E^{\mathcal{B}}_{\varphi(g)}\big{(}\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\big{)}=E^{\mathcal{B}}_{\varphi(g)}\big{(}\Lambda^{\mathcal{B}}(\Phi_{T}(\widehat{j}^{\,{\mathcal{A}}}_{g}(a)))\big{)},$
and the assertion follows from Proposition 3.8 (with $\Phi=\Phi_{T}$).
We will now discuss whether the converse of these two assertions hold. We will first show that it does in the full case. To obtain a reduced version, we will have to assume that the kernel of $\varphi$ is amenable. Such a condition is known to be necessary and sufficient when considering the extendibility of the homomorphism $\varphi:G\to H$ to a *-homomorphism from $C_{r}^{*}(G)$ into $C_{r}^{*}(H)$ (cf. [11, Corollary 8.C.15]). Note here that a completely positive map from $C_{r}^{*}(G)$ into $C_{r}^{*}(H)$ which extends $\varphi$ has to be a *-homomorphism, so one gets from the result cited above that $\varphi$ extends to a completely positive linear map from $C_{r}^{*}(G)$ into $C_{r}^{*}(H)$ if and only if ${\rm ker}(\varphi)$ is amenable.
Theorem 3.10.
Let $T=(T_{g})_{g\in G}$ be an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. Then $T$ is positive definite if and only if $T$ is full and the associated map $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ is completely positive. In such a case, we have $\|\Phi_{T}\|=\|T_{e}\|$.
Proof.
In view of our preliminary discussion, in order to prove the first statement, we only need to show the forward implication. So assume that $T$ is positive definite. Let $\psi_{T}:{\mathcal{A}}\to C^{*}({\mathcal{B}})$ be the map given by
$\psi_{T}(a)=\widehat{j}^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))$
whenever $a\in A_{g}$ for $g\in G$, and let $\Pi$ be a faithful *-representation of $C^{*}({\mathcal{B}})$ on some Hilbert space ${\mathcal{H}}$.
As a first step, we will show that $\phi:=\Pi\circ\psi_{T}:{\mathcal{A}}\to B({\mathcal{H}})$ is completely positive in the sense of Buss, Ferraro and Sehnem, cf. [18, Appendix]. Let $g_{1},\ldots,g_{n}\in G$, set $\mathbf{g}:=(g_{1},\ldots,g_{n})\in G^{n}$ and let $\phi^{\mathbf{g}}:M_{\mathbf{g}}({\mathcal{A}})\to M_{n}(B({\mathcal{H}}))$ be the linear map defined by
$\phi^{\mathbf{g}}([r_{ij}])=[\phi(r_{ij})]=\big{[}\Pi(\psi_{T}(r_{ij}))\big{]}\,\text{ for each }[r_{ij}]\in M_{\mathbf{g}}({\mathcal{A}}).$
To show that $\phi$ is completely positive amounts to check that $\phi^{\mathbf{g}}$ is a positive map for every $\mathbf{g}\in\cup_{n\in{\mathbb{N}}}\,G^{n}$. It clearly suffices to show that the linear map $\psi^{\mathbf{g}}_{T}:M_{\mathbf{g}}({\mathcal{A}})\to M_{n}(C^{*}({\mathcal{B}}))$, given by
$\psi^{\mathbf{g}}_{T}([r_{ij}])=\big{[}\psi_{T}(r_{ij})\big{]}=\Big{[}\,\widehat{j}^{\mathcal{B}}_{\varphi(g_{i}^{-1}g_{j})}(T_{g_{i}^{-1}g_{j}}(r_{ij}))\Big{]}\,\text{ for each }R=[r_{ij}]\in M_{\mathbf{g}}({\mathcal{A}}),$
is a positive map for every $\mathbf{g}\in\cup_{n\in{\mathbb{N}}}\,G^{n}$. So let $\mathbf{g}=(g_{1},\ldots,g_{n})\in G^{n}$ and consider $a_{i}\in A_{g_{i}}$ for $i=1,\ldots,n$. Since $T$ is positive definite, the matrix $\big{[}T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\big{]}$ is positive in $M_{\varphi(\mathbf{g})}({\mathcal{B}})$. Now, it is easy to check that if $\mathbf{h}=(h_{1},\ldots,h_{n})\in H^{n}$, then the map
$\big{[}s_{ij}\big{]}\in M_{\mathbf{h}}({\mathcal{B}})\mapsto\Big{[}\,\widehat{j}^{\mathcal{B}}_{h_{i}^{-1}h_{j}}(s_{ij})\Big{]}\in M_{n}(C^{*}({\mathcal{B}}))$
is a *-homomorphism. Hence we get that
$\psi^{\mathbf{g}}_{T}\big{(}[a_{i}^{*}a_{j}]\big{)}=\Big{[}\,\widehat{j}^{\mathcal{B}}_{\varphi(g_{i})^{-1}\varphi(g_{j})}(T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j}))\Big{]}\geq 0\,.$
It then follows that $\psi_{T}^{\mathbf{g}}([r_{ij}])\geq 0$ for every $R=[r_{ij}]\in M_{\mathbf{g}}({\mathcal{A}})^{+}$, hence that $\psi_{T}^{\mathbf{g}}$ is a positive map.
Indeed, assume $R=C^{*}C$ for some $C=[c_{ij}]\in M_{\mathbf{g}}({\mathcal{A}})$. Then $R=\sum_{k=1}^{n}\big{[}c_{ki}^{*}c_{kj}\big{]},$ so applying what we have just shown gives that
$\psi^{\mathbf{g}}_{T}(R)=\sum_{k=1}^{n}\psi^{\mathbf{g}}_{T}\big{(}\big{[}c_{ki}^{*}c_{kj}\big{]}\big{)}$
is positive in $M_{n}(C^{*}({\mathcal{B}}))$, being a finite sum of positive elements.
This shows that $\phi=\Pi\circ\psi_{T}:{\mathcal{A}}\to B({\mathcal{H}})$ is completely positive, as desired.
By [18, Corollary A.8], there exists a completely positive map $\tilde{\phi}:C^{*}({\mathcal{A}})\to B({\mathcal{H}})$ such that
$\tilde{\phi}\big{(}\widehat{j}^{\mathcal{A}}_{g}(a)\big{)}=\phi(a)\quad\text{for every }g\in G\text{ and }a\in A_{g}.$
We then get that
$\tilde{\phi}(\widehat{j}^{\mathcal{A}}_{g}(a))=\phi(a)=\Pi(\psi_{T}(a))=\Pi\big{(}\,\widehat{j}^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\big{)}\quad\text{forevery }g\in G\text{ and }a\in A_{g}.$
Thus, $\tilde{\phi}\big{(}C^{*}({\mathcal{A}})\big{)}\subseteq\Pi(C^{*}({\mathcal{B}}))$. Since $\Pi$ is injective, we may regard it as a *-isomorphism onto its range $\Pi(C^{*}({\mathcal{B}}))$ and define $\phi^{\prime}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ by $\phi^{\prime}=\Pi^{-1}\circ\tilde{\phi}$. Then $\phi^{\prime}$ is completely positive and satisfies
$\phi^{\prime}\big{(}\widehat{j}^{\mathcal{A}}_{g}(a)\big{)}=\widehat{j}^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\quad\text{for every }g\in G\text{ and }a\in A_{g}.$
Thus, $\phi^{\prime}$ extends $\phi_{T}$, so $T$ is full, and $\Phi_{T}=\phi^{\prime}$ is completely positive, as we wanted to show.
Finally, let $\{u_{i}\}_{i\in I}$ be an approximate unit for $A_{e}$. Then $\big{\{}\widehat{j}_{e}^{\mathcal{A}}(u_{i})\big{\}}_{i\in I}$ is an approximate unit for $C^{*}({\mathcal{A}})$. So if $T$ is positive definite, then $T_{e}$ and $\Phi_{T}$ are completely positive, and thus
$\|\Phi_{T}\|=\lim_{i}\|\Phi_{T}(\,\widehat{j}_{e}^{\mathcal{A}}(u_{i}))\|_{\rm u}=\lim_{i}\|\widehat{j}_{e}^{\mathcal{A}}(T_{e}(u_{i}))\|_{\rm u}=\lim_{i}\|T_{e}(u_{i})\|=\|T_{e}\|.$
$\square$
Note that if ${\mathcal{A}}$ and ${\mathcal{B}}$ are amenable in the sense of Exel [26, 27], i.e., their full and reduced cross-sectional $C^{*}$-algebras are canonically *-isomorphic, respectively, then we get as a corollary to Theorem 3.10 that $T$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map if and only if $T$ is reduced and $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ is completely positive. When ${\mathcal{A}}$ or ${\mathcal{B}}$ are not amenable, this equivalence is also true under the assumption that ${\rm ker}(\varphi)$ is an amenable subgroup of $G$. In order to show this, we will first generalize [18, Proposition 4.8].
Proposition 3.11.
Assume that the kernel of $\varphi:G\to H$ is amenable and let $\pi=(\pi_{h})_{h\in H}$ be a representation of ${\mathcal{B}}$ on a Hilbert space ${\mathcal{H}}$ such that $\pi_{e}$ is faithful. Suppose that $\psi:C^{*}({\mathcal{A}})\to B({\mathcal{H}})$ is a completely positive linear map such that for all $g\in G$
$\psi(\,\widehat{j}^{\mathcal{A}}_{g}(A_{g}))\subseteq\,\pi_{\varphi(g)}(B_{\varphi(g)}).$
Then there is a completely positive linear map $\psi^{\prime}:C^{*}_{r}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ such that for all $g\in G$
$\psi^{\prime}\big{(}\lambda^{\mathcal{A}}_{g}(A_{g})\big{)}\subseteq\,\lambda^{\mathcal{B}}_{\varphi(g)}(B_{\varphi(g)})$ (3)
and
$\big{(}\pi_{\varphi(g)}\circ(\lambda^{\mathcal{B}}_{\varphi(g)})^{-1}\circ\psi^{\prime}\circ{\lambda^{\mathcal{A}}_{g}}\big{)}_{|A_{g}}=\big{(}\psi\circ\widehat{j}^{\mathcal{A}}_{g}\big{)}_{|A_{g}}.$ (4)
Proof.
By Stinespring’s dilation theorem for completely positive maps (see e.g. [46, Theorem 5.6]), there exist a Hilbert space ${\mathcal{K}}$, a *-homomorphism $\widehat{\psi}:C^{*}({\mathcal{A}})\to B({\mathcal{K}})$ and a bounded operator $V:{\mathcal{H}}\to{\mathcal{K}}$ such that $\psi(x)=V^{*}\,\widehat{\psi}(x)\,V$ for all $x\in C^{*}({\mathcal{A}})$. Let $\widehat{\pi}=(\widehat{\pi}_{g})_{g\in G}$ denote the representation of ${\mathcal{A}}$ on ${\mathcal{K}}$ associated to $\widehat{\psi}$, so that
$\widehat{\psi}(\widehat{j}_{g}^{\mathcal{A}}(a))=\widehat{\pi}_{g}(a)\quad\text{ for all }g\in G\text{~and }~a\in A_{g}.$
By Fell’s absorption principle for Fell bundles [27, Proposition 18.4], the integrated forms of the representations $\widehat{\pi}\otimes\lambda^{G}=(\widehat{\pi}_{g}\otimes\lambda^{G}_{g})_{g\in G}$ and $\pi\otimes\lambda^{H}=(\pi_{h}\otimes\lambda^{H}_{h})_{h\in H}$ factor through $C^{*}_{r}({\mathcal{A}})$ and $C^{*}_{r}({\mathcal{B}})$, respectively. Let $\psi_{\mathcal{A}}:C^{*}_{r}({\mathcal{A}})\to B({\mathcal{K}}\otimes\ell^{2}(G))$ and $\psi_{\mathcal{B}}:C^{*}_{r}({\mathcal{B}})\to B({\mathcal{H}}\otimes\ell^{2}(H))$ denote the induced *-homomorphisms. In fact, using [27, Corollary 18.5], we may regard $\Psi_{\mathcal{A}}$ as a map from $C^{*}_{r}({\mathcal{A}})$ into $B({\mathcal{K}})\otimes_{\rm min}C_{r}^{*}(G)$ and $\Psi_{\mathcal{B}}$ as a map from $C^{*}_{r}({\mathcal{B}})$ into $B({\mathcal{H}})\otimes_{\rm min}C_{r}^{*}(H)$, satisfying that
$\psi_{\mathcal{A}}(\lambda_{g}^{\mathcal{A}}(a))=\widehat{\pi}_{g}(a)\otimes\lambda_{g}^{G}\text{~and }\psi_{\mathcal{B}}(\lambda_{h}^{\mathcal{B}}(b))=\pi_{h}(b)\otimes\lambda_{h}^{H}$
for all $g\in G$, $a\in A_{g}$, $h\in H$ and $b\in B_{h}$. Note that $\psi_{\mathcal{B}}$ is faithful since $\pi_{e}$ is assumed to be so, cf. [27, Proposition 18.4].
Now, as shown in [11, Corollary 8.C.15], $\varphi$ extends to a *-homomorphism $\varphi_{*,r}:C_{r}^{*}(G)\to C_{r}^{*}(H)$ satisfying that $\varphi_{*,r}(\lambda_{g}^{G})=\lambda^{H}_{\varphi(g)}$ for all $g\in G$.
Let $\widetilde{V}:B({\mathcal{K}})\to B({\mathcal{H}})$ denote the completely positive map given by $\widetilde{V}(T)=V^{*}TV$ for all $T\in B({\mathcal{K}})$. Using [14, Theorem 3.5.3], we may define a completely positive map $\Phi:C_{r}^{*}({\mathcal{A}})\to B({\mathcal{H}})\otimes_{\rm min}C_{r}^{*}(H)$ by
$\Phi:=(\widetilde{V}\otimes_{\rm min}\varphi_{*,r})\circ\Psi_{\mathcal{A}}.$
Let $g\in G$ and $a\in A_{g}$. Then we have
$~\Phi(\lambda^{\mathcal{A}}_{g}(a))$ $=(\widetilde{V}\otimes_{\rm min}\varphi_{*,r})(\psi_{\mathcal{A}}(\lambda^{\mathcal{A}}_{g}(a)))$
$=(\widetilde{V}\otimes_{\rm min}\varphi_{*,r})(\widehat{\pi}_{g}(a)\otimes\lambda^{G}_{g})$
$=V^{*}\widehat{\pi}_{g}(a)V\otimes\lambda^{H}_{\varphi(g)}$
$=V^{*}\widehat{\psi}(\widehat{j}^{\mathcal{A}}_{g}(a))V\otimes\lambda^{H}_{\varphi(g)}$
$=\psi(\,\widehat{j}^{\mathcal{A}}_{g}(a))\otimes\lambda^{H}_{\varphi(g)}.$
By assumption, we have that $\psi(\,\widehat{j}^{\mathcal{A}}_{g}(a))\in\pi_{\varphi(g)}(B_{\varphi(g)})$. Thus there is some $b\in B_{\varphi(g)}$ such that $\psi(\,\widehat{j}^{\mathcal{A}}_{g}(a))=\pi_{\varphi(g)}(b)$. Hence
$\Phi(\lambda^{\mathcal{A}}_{g}(a))=\pi_{\varphi(g)}(b)\otimes\lambda^{H}_{\varphi(g)}=\psi_{\mathcal{B}}(\lambda^{\mathcal{B}}_{\varphi(g)}(b)).$
By a density and continuity argument, we obtain that $\Phi(C_{r}^{*}({\mathcal{A}}))\subseteq\psi_{\mathcal{B}}(C_{r}^{*}({\mathcal{B}}))$.
Since $\psi_{\mathcal{B}}$ is faithful, we can define $\psi^{\prime}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ by $\psi^{\prime}=(\psi_{\mathcal{B}})^{-1}\circ\Phi$, which then satisfies $\psi^{\prime}(\lambda^{\mathcal{A}}_{g}(A_{g}))\subseteq\lambda^{\mathcal{B}}_{\varphi(g)}(B_{\varphi(g)})$, i.e., (3) holds. It is now straightforward to check that $\psi^{\prime}$ also satisfies (4). $\square$
Theorem 3.12.
Let $T=(T_{g})_{g\in G}$ be an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map, and assume that the kernel of $\varphi$ is amenable. Then $T$ is positive definite if and only if $T$ is reduced and the associated map $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ is completely positive, in which case we have $\|M_{T}\|~=\|T_{e}\|$.
Proof.
As we have seen previously, the backward implication is always true. So assume that $T=(T_{g})_{g\in G}$ is positive definite. By Theorem 3.10 we know that $\phi_{T}$ extends to a completely positive map $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ satisfying that $\Phi_{T}\big{(}\widehat{j}^{\mathcal{A}}_{g}(a)\big{)}=\widehat{j}^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))$ for all $g\in G$ and $a\in A_{g}$. Let $\pi=(\pi_{h})_{h\in H}$ be a *-representation of ${\mathcal{B}}$ in $B({\mathcal{H}})$ for some Hilbert space ${\mathcal{H}}$ such that its integrated form $\Pi:C^{*}({\mathcal{B}})\to B({\mathcal{H}})$ is faithful. Note that $\pi_{e}=\Pi\circ\widehat{j}^{\,{\mathcal{B}}}_{e}$ is faithful too. Then the linear map $\widetilde{\phi}:=\Pi\circ\Phi_{T}:C^{*}({\mathcal{A}})\to B({\mathcal{H}})$ is completely positive and satisfies that
$\widetilde{\phi}\Big{(}\widehat{j}^{\mathcal{A}}_{g}(a)\Big{)}=\Pi\big{(}\Phi_{T}(\widehat{j}^{\mathcal{A}}_{g}(a))\big{)}=\Pi\big{(}\widehat{j}^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\big{)}=\pi_{\varphi(g)}(T_{g}(a))$
for all $g\in G$ and $a\in A_{g}$. Now, using Proposition 3.11, we get that there exists a completely positive map $\phi^{\prime}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ such that $\phi^{\prime}\big{(}\lambda^{\mathcal{A}}_{g}(A_{g})\big{)}\subseteq\lambda^{\mathcal{B}}_{\varphi(g)}(B_{\varphi(g)})$ and
$\big{(}\pi_{\varphi(g)}\circ(\lambda^{\mathcal{B}}_{\varphi(g)})^{-1}\circ\phi^{\prime}\circ{\lambda^{\mathcal{A}}_{g}}\big{)}_{|A_{g}}=\big{(}\widetilde{\phi}\circ\widehat{j}^{\mathcal{A}}_{g}\big{)}_{|A_{g}}$
for each $g\in G$. This means that
$\pi_{\varphi(g)}\big{(}(\lambda^{\mathcal{B}}_{\varphi(g)})^{-1}(\phi^{\prime}(\lambda^{\mathcal{A}}_{g}(a)))\big{)}=\pi_{\varphi(g)}(T_{g}(a))$
for $g\in G$ and $a\in A_{g}$. Since $\pi_{\varphi(g)}$ is injective on $B_{\varphi(g)}$, this implies that
$\phi^{\prime}(\lambda^{\mathcal{A}}_{g}(a))=\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\quad\text{for every }g\in G\text{ and }a~\in A_{g}.$
This shows that $\phi^{\prime}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ is a bounded extension of $\phi_{T}$. Hence $T$ is reduced, and $M_{T}=\phi^{\prime}$ is completely positive, as desired.
Finally, if $T$ is positive definite, then one may proceed in a similar way as in the full case to show that $\|M_{T}\|=\|T_{e}\|$. $\square$
Remark 3.13.
Assume ${\rm ker}(\varphi)$ is not amenable and $T=(T_{g})_{g\in G}$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. Then it may still happen that $T$ is reduced and $M_{T}$ is completely positive, cf. Example 4.1. Hence, in such a setting, it would be interesting to find some additional condition(s) ensuring that $T$ is reduced with $M_{T}$ completely positive.
Remark 3.14.
One could say that a completely positive linear map $\Psi:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ is $\varphi$-diagonal whenever
$\Psi(\,\widehat{j}^{\,{\mathcal{A}}}_{g}(A_{g}))\subseteq\,\,\widehat{j}^{\,{\mathcal{B}}}_{\varphi(g)}(B_{\varphi(g)})$
for all $g\in G$. Note that if $T$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map, then $\Phi_{T}$ is $\varphi$-diagonal. For a general completely positive linear map $\Phi:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$, letting $T^{\Phi}$ denote the positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map associated to it in Proposition 3.7, we get from Theorem 3.10 that the map $\Phi_{T^{\Phi}}$ is a $\varphi$-diagonal completely positive linear map (which could be called the $\varphi$-diagonalization of $\Phi$). Hence the map $\Phi\mapsto\Phi_{T^{\Phi}}$ is a projection from the space CP($C^{*}({\mathcal{A}}),C^{*}({\mathcal{B}})$) of completely positive linear maps onto the subspace of $\varphi$-diagonal maps. If the kernel of $\varphi$ is amenable, using Proposition 3.8 and Theorem 3.12, one obtains a similar projection map on $CP(C_{r}^{*}({\mathcal{A}}),C_{r}^{*}({\mathcal{B}}))$.
Remark 3.15.
The family of ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle maps forms a vector space $L({\mathcal{A}},\varphi,{\mathcal{B}})$ with respect to pointwise addition and scalar multiplication. Set
$B({\mathcal{A}},\varphi,{\mathcal{B}}):={\rm span}\{T\in L({\mathcal{A}},\varphi,{\mathcal{B}}):T\text{ is positive definite}\}.$
Then any $S\in B({\mathcal{A}},\varphi,{\mathcal{B}})$ is a full ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map, which is also reduced if ${\rm ker}(\varphi)$ is assumed to be amenable. This follows readily from Theorems 3.10 and 3.12 after noticing that the map $T\mapsto\phi_{T}$ is linear. We thus get a linear map $T\mapsto\Phi_{T}$ from $B({\mathcal{A}},\varphi,{\mathcal{B}})$ into the space of completely bounded linear maps CB($C^{*}({\mathcal{A}}),C^{*}({\mathcal{B}})$), and a similar one into CB($C_{r}^{*}({\mathcal{A}}),C_{r}^{*}({\mathcal{B}})$) when ${\rm ker}(\varphi)$ is amenable. It is easy to see that the space $B({\mathcal{A}}):=B({\mathcal{A}},{\rm id}_{G},{\mathcal{A}})$ becomes a unital algebra w.r.t. natural composition, which could be called the Fourier-Stieltjes algebra of ${\mathcal{A}}$, in analogy with the case where ${\mathcal{A}}$ is the Fell bundle associated to a unital discrete twisted $C^{*}$-dynamical system, cf. [9].
When $G=H$ and $\varphi={\rm id}_{G}$, we may combine Theorem 3.10 and Theorem 3.12 to obtain the following.
Corollary 3.16.
Assume ${\mathcal{A}}$ and ${\mathcal{B}}$ are both Fell bundles over $G$ and let $T=(T_{g})_{g\in G}$ be an ${\mathcal{A}}$$-{\mathcal{B}}$ bundle map. Then we have that $T$ is positive definite if and only if $T$ is full and the associated map $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ is completely positive, if and only if $T$ is reduced and the associated map $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ is completely positive. When this happens, we have $\|\Phi_{T}\|=\|M_{T}\|~=\|T_{e}\|$.
3.3 Categorical considerations
There is a category ${\tt Fell}$-${\tt PD}$ whose objects are Fell bundles over discrete groups, and an arrow from ${\mathcal{A}}=(A_{g})_{g\in G}$ to ${\mathcal{B}}=(B_{h})_{h\in H}$ is defined as a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map for some homomorphism $\varphi:G\to H$, with composition of arrows defined in the natural way. More formally, we define an arrow from ${\mathcal{A}}=(A_{g})_{g\in G}$ to ${\mathcal{B}}=(B_{h})_{h\in H}$ to be a pair $(\varphi,T)$, where $\varphi:G\to H$ is a homomorphism and $T$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. Composition of an arrow $(\varphi,T)$ from ${\mathcal{A}}$ to ${\mathcal{B}}$ with an arrow $(\psi,S)$ from ${\mathcal{B}}$ to ${\mathcal{C}}=(C_{k})_{k\in K}$ is defined by
$(\psi,S)\circ(\varphi,T)=(\psi\varphi,R)$
where $R$ is the ${\mathcal{A}}$-$\psi\varphi$-${\mathcal{C}}$ bundle map given by $R:=(S_{\varphi(g)}T_{g})_{g\in G}$. Note here that $R$ is positive definite, hence that $(\psi\varphi,R)$ is an arrow in ${\tt Fell}$-${\tt PD}$. Indeed, as
$(\Phi_{S}\circ\Phi_{T})\Big{(}\widehat{j}^{\mathcal{A}}_{g}(b)\Big{)}=\Phi_{S}\Big{(}\widehat{j}^{\mathcal{B}}_{\varphi(g)}(T_{g}(b))\Big{)}=\widehat{j}^{\mathcal{C}}_{\psi(\varphi(g))}(S_{\varphi(g)}(T_{g}(b)))=\widehat{j}^{\mathcal{C}}_{(\psi\varphi)(g)}(R_{g}(b)))$
for all $g\in G\text{ and}~b\in B_{g}$, we get that $\Phi_{S}\circ\Phi_{T}$ is a completely positive linear map from $C^{*}({\mathcal{A}})$ into $C^{*}({\mathcal{C}})$ which extends $\phi_{R}$, and it follows from Theorem 3.10 that $R$ is positive definite.
For an object ${\mathcal{A}}$ in ${\tt Fell}$-${\tt PD}$, the identity arrow is defined as $({\rm id}_{G},{\rm id}_{\mathcal{A}})$ where ${\rm id}_{G}$ is the identity map on $G$ and ${\rm id}_{\mathcal{A}}=(({\rm id}_{\mathcal{A}})_{g})_{g\in G}$ with $({\rm id}_{\mathcal{A}})_{g}(a):=a$ for all $g\in G$ and $a\in A_{g}$. It is now straightforward to check that ${\tt Fell}$-${\tt PD}$ becomes a category.
The main point with introducing ${\tt Fell}$-${\tt PD}$ is that we get a functor $\mathcal{F}$ from the category ${\tt Fell}$-${\tt PD}$ to the category ${\tt C^{*}CP}$ whose objects are $C^{*}$-algebras and arrows are completely positive linear maps, by setting
$\mathcal{F}({\mathcal{A}}):=C^{*}({\mathcal{A}})\,\text{ and }\,\mathcal{F}(\varphi,T):=\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$
for an object ${\mathcal{A}}$ in ${\tt Fell}$-${\tt PD}$ and an arrow $(\varphi,T)$ from ${\mathcal{A}}$ to ${\mathcal{B}}$ in ${\tt Fell}$-${\tt PD}$.
The category ${\tt Fell}$-${\tt PD}$ contains a subcategory ${\tt Fell}$ with the same objects, but where an arrow $(\varphi,T)$ from ${\mathcal{A}}$ to ${\mathcal{B}}$ is required to be a morphism according to Definition 3.1. As $\Phi_{T}=\mathcal{F}(\varphi,T)$ is a *-homomorphism from $C^{*}({\mathcal{A}})$ into $C^{*}({\mathcal{B}})$ whenever $(\varphi,T)$ is a morphism, we may restrict $\mathcal{F}$ to ${\tt Fell}$ and get a functor from ${\tt Fell}$ into the category ${\tt C^{*}HOM}$ of $C^{*}$-algebras having *-homomorphisms as arrows.
We may also introduce the subcategory ${\tt Fell}$-${\tt PD}_{\tt red}$ of ${\tt Fell}$-${\tt PD}$ having the same objects but whose arrows involve only group homomorphisms with amenable kernels. The fact that composition of arrows in this category makes sense is due to following result: if $\varphi:G\to H$ and $\psi:H\to K$ are group homomorphisms such that ${\rm ker}(\varphi)$ and ${\rm ker}(\psi)$ are amenable, then ${\rm ker}(\psi\varphi)$ is amenable. This can be shown by first observing that ${\rm ker}(\varphi)$ is a normal subgroup of ${\rm ker}(\psi\varphi)$ such that the quotient group ${\rm ker}(\psi\varphi)/{\rm ker}(\varphi)$ is isomorphic to a subgroup of ${\rm ker}(\psi)$, from which the amenability of ${\rm ker}(\psi\varphi)$ then follows by using that amenability is preserved when passing to subgroups and closed under extensions.
We then get a functor $\mathcal{F}_{\rm red}$ from ${\tt Fell}$-${\tt PD}_{\tt red}$ to ${\tt C^{*}CP}$ by setting
$\mathcal{F}_{\rm red}({\mathcal{A}}):=C_{r}^{*}({\mathcal{A}})\text{ and }\mathcal{F}_{\rm red}(\varphi,T):=M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$
for an object ${\mathcal{A}}$ in ${\tt Fell}$-${\tt PD}_{\tt red}$ and an arrow $(\varphi,T)$ from ${\mathcal{A}}$ to ${\mathcal{B}}$ in ${\tt Fell}$-${\tt PD}_{\tt red}$.
The category ${\tt Fell}$-${\tt PD}_{\tt red}$ contains an obvious subcategory ${\tt Fell}_{\tt red}$ having the same objects, but whose arrows are morphisms $(\varphi,T):{\mathcal{A}}\to{\mathcal{B}}$ with ${\rm ker}(\varphi)$ is amenable. By restricting $\mathcal{F}_{\rm red}$ to ${\tt Fell}_{\tt red}$ we get a functor from ${\tt Fell}_{\tt red}$ into ${\tt C^{*}HOM}$.
Sometimes, the subcategory ${\tt C^{*}CCP}$ of ${\tt C^{*}CP}$ whose arrows consist of contractive completely positive maps is the one of interest. It follows from Theorems 3.10 and 3.12 that we obtain functors into this category by restricting $\mathcal{F}$ and $\mathcal{F}_{\rm red}$ to the respective subcategories of ${\tt Fell}$-${\tt PD}$ and ${\tt Fell}$-${\tt PD}_{\tt red}$ whose arrows are such that $T_{e}$ is contractive.
4 Some examples
Example 4.1.
Let’s first consider the simple case where ${\mathcal{A}}=({\mathbb{C}}\times\{g\})_{g\in G}$ is the group bundle associated to $G$, =×${\mathcal{B}}={\mathbb{C}}\times\{e\}$ is the trivial bundle over $\{e\}$ and $\varphi_{0}:G\to\{e\}$ is the trivial homomorphism. Then one readily sees that an ${\mathcal{A}}$-$\varphi_{0}$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ is of the form $T_{g}(z,g)=(\phi(g)z,e)$ for some (necessarily bounded) function $\phi:G\to{\mathbb{C}}$, and that $T$ is positive definite if and only if $\phi$ is positive definite. In this case, we have $C^{*}({\mathcal{A}})=C^{*}(G)$, $C^{*}({\mathcal{B}})={\mathbb{C}}$, and the map $T\mapsto\Phi_{T}$ corresponds to the usual correspondence between positive definite functions on $G$ and positive linear functionals on $C^{*}(G)$.
Moreover, we have $C_{r}^{*}({\mathcal{A}})=C_{r}^{*}(G)$, $C_{r}^{*}({\mathcal{B}})={\mathbb{C}}$. Thinking of positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathcal{B}}$ bundle maps as positive definite functions on $G$, it is clear that $M_{T}$ is defined exactly when $T$ is a positive definite function associated with a unitary representation of $G$ weakly contained in $\lambda^{G}$, in which case $M_{T}$ is the corresponding positive linear functional on $C_{r}^{*}(G)$. Note that this is true regardless of whether $G={\rm ker}(\varphi_{0})$ is amenable or not.
Example 4.2.
Next, if both ${\mathcal{A}}$ and ${\mathcal{B}}$ are Fell bundles over the trivial group $\{e\}$ with fibers $A$ and $B$, respectively, then a positive definite ${\mathcal{A}}$$-{\mathcal{B}}$ bundle map $T$ corresponds to a completely positive map from $A$ into $B$. Of course, in this case, we have $C^{*}({\mathcal{A}})=A$ and $C^{*}({\mathcal{B}})=B$.
Example 4.3.
We now look at a more general situation, covering both the previous examples. Let ${\mathcal{A}}=(A_{g})_{g\in G}$ be a Fell bundle and $B$ be a $C^{*}$-algebra. Let ${\mathcal{B}}$ be the Fell bundle over the trivial group $\{e\}$ with fiber $B$ over $e$, so $B=C^{*}({\mathcal{B}})$, and let $\varphi_{0}:G\to\{e\}$ be the trivial homomorphism. Then a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ is a family of linear maps $T_{g}:A_{g}\to B$ such that the matrix $\big{[}T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\big{]}$ is positive in $M_{n}(B)$ for every $g_{1},\ldots,g_{n}\in G$ and $a_{i}\in A_{g_{i}},i=1,\ldots,n$.
In particular, choosing $B={\mathbb{C}}$, we get that a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathbb{C}}$ bundle map $\omega=(\omega_{g})_{g\in G}$ is a family of linear functionals $\omega_{g}:A_{g}\to{\mathbb{C}}$ such that the matrix $\big{[}\omega_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\big{]}$ is positive in $M_{n}({\mathbb{C}})$ for every $g_{1},\ldots,g_{n}\in G$ and $a_{i}\in A_{g_{i}},i=1,\ldots,n$. This is easily seen to be equivalent to requiring that
[…]
for every $g_{1},\ldots,g_{n}\in G$ and $a_{i}\in A_{g_{i}},i=1,\ldots,n$. Such a bundle map $\omega=(\omega_{g})_{g\in G}$ is essentially the same as what Fell and Doran call a functional on ${\mathcal{A}}$ of positive type, cf. [29, Chapter VIII, 20.2]. We refer to [4] for a recent study of such maps. In accordance with Theorem 3.10, every such a functional $\omega=(\omega_{g})_{g\in G}$ on ${\mathcal{A}}$ gives rise to a positive linear functional $\Phi_{\omega}$ on $C^{*}({\mathcal{A}})$ satisfying that $\Phi_{\omega}(\,\widehat{j}_{g}^{\mathcal{A}}(a))=\omega_{g}(a)$ for $g\in G$ and $a\in A_{g}$. Conversely, each positive linear functional $\Psi$ on $C^{*}({\mathcal{A}})$ gives rise to a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathbb{C}}$ bundle map $\omega^{\Psi}=(\omega^{\Psi}_{g})_{g\in G}$ given by $\omega_{g}^{\Psi}(a)=\Psi(\,\widehat{j}_{g}^{\mathcal{A}}(a))$ for $g\in G$ and $a\in A_{g}$, which satisfies that $\Psi=\Phi_{\omega^{\Psi}}$. It follows that the map $\omega\mapsto\Phi_{\omega}$ is a bijection between the cone of positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathbb{C}}$ bundle maps and the cone of positive linear functionals on $C^{*}({\mathcal{A}})$.
In another direction, let now ${\mathcal{A}}=({\mathbb{C}}\times\{g\})_{g\in G}$ be the group bundle over $G$ and assume $B=B({\mathcal{H}})$ for some Hilbert space ${\mathcal{H}}$. Then a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathcal{B}}$ bundle map amounts to a map $T:G\to B({\mathcal{H}})$ which is completely positive definite in the sense of Paulsen [56, Chapter 4, p. 51] (some other authors just say positive definite, e.g. [13, Section 6, p. 115] or [53, Section 8.3]), i.e., it satisfies that the matrix $[T(g_{i}^{-1}g_{j})]$ is positive in $M_{n}(B({\mathcal{H}}))$ for every $g_{1},\ldots,g_{n}\in G$. As mentioned by Paulsen on p. 54, there is a one-to-one correspondence between such completely positive definite maps from $G$ into $B({\mathcal{H}})$ and completely positive maps from $C^{*}(G)$ into $B({\mathcal{H}})$. This fits with Theorem 3.10.
Example 4.4.
Consider a morphism $(\varphi,T)$ from a Fell bundle ${\mathcal{A}}=(A_{g})_{g\in G}$ to a Fell bundle ${\mathcal{B}}=(B_{h})_{h\in H}$. Then $T$ is positive definite. Indeed, let $g_{1},\ldots,g_{n}\in G$, $a_{i}\in A_{g_{i}}$ and $b_{i}\in B_{\varphi(g_{i})}$ for $i=1,\ldots,n$. Then
$\sum_{i,j=1}^{n}b_{i}\,T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\,b_{j}^{*}$ $=\sum_{i,j=1}^{n}b_{i}\,T_{g_{i}^{-1}}(a_{i}^{*})T_{g_{j}}(a_{j})\,b_{j}^{*}=\sum_{i,j=1}^{n}b_{i}\,T_{g_{i}}(a_{i})^{*}T_{g_{j}}(a_{j})\,b_{j}^{*}$
$=\Big{(}\sum_{i=1}^{n}T_{g_{i}}(a_{i})b_{i}^{*}\Big{)}^{*}\Big{(}\sum_{j=1}^{n}T_{g_{j}}(a_{j})\,b_{j}^{*}\Big{)}\in(B_{e})^{+}.$
We can now conclude from Theorem 3.10 that there is a unique completely positive linear map $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ satisfying that
$\Phi_{T}(\widehat{j}^{\,{\mathcal{A}}}_{g}(a))=\widehat{j}^{\,{\mathcal{B}}}_{\varphi(g)}(T_{g}(a))$
for all $g\in G$ and $a\in A_{g}$. As $T$ is multiplicative, one readily checks that $\phi_{T}:C_{c}({\mathcal{A}})\to C_{c}({\mathcal{B}})$ is multiplicative. Since its extension $\Phi_{T}$ is bounded, this implies that $\Phi_{T}$ is multiplicative, hence that $\Phi_{T}$ is a *-homomorphism.
Similarly, if one assumes that ${\rm ker}(\varphi)$ is amenable, we get from Theorem 3.12 that there is a unique *-homomorphism $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ satisfying that
$M_{T}(\lambda^{\mathcal{A}}_{g}(a))=\lambda^{\mathcal{B}}_{g}(T_{g}(a))$
for all $g\in G$ and $a\in A_{g}$.
These observations generalize [27, Propositions 21.2 and 21.3], where the case $G=H$ and $\varphi={\rm id}_{G}$ is considered.
Example 4.5.
Assume ${\mathcal{B}}=(B_{g})_{g\in G}$ is a Fell bundle. Let $\phi:G\to{\mathbb{C}}$ be positive definite and let $T^{\phi}=(T^{\phi}_{g})_{g\in G}$ be the ${\mathcal{B}}$-bundle map given by
$T^{\phi}_{g}(b)=\phi(g)b$
for every $g\in G$ and $b\in B_{g}$. Then $T^{\phi}$ is positive definite. Indeed, for $g_{1},\ldots,g_{n}\in G$ and $b_{i},c_{i}\in B_{g_{i}}$, $i=1,\ldots,n$, we have
$\sum_{i,j=1}^{n}c_{i}\,T^{\phi}_{g_{i}^{-1}g_{j}}(b_{i}^{*}b_{j})\,c_{j}^{*}=~\sum_{i,j=1}^{n}c_{i}\,\phi(g_{i}^{-1}g_{j})b_{i}^{*}b_{j}\,c_{j}^{*}=~\sum_{i,j=1}^{n}\phi(g_{i}^{-1}g_{j})\,a_{ij}$
where $a_{ij}:=c_{i}b_{i}^{*}b_{j}c_{j}^{*}\in B_{e}$. Now, the matrix $[\phi(g_{i}^{-1}g_{j})]$ is positive in $M_{n}({\mathbb{C}})$ if $\phi$ is positive definite, and $[a_{ij}]$ is clearly positive in $M_{n}(B_{e})$. Hence, the matrix $[\phi(g_{i}^{-1}g_{j})a_{ij}]$ is then positive in $M_{n}(B_{e})$ (cf. [47, Lemma 3.1]) and the claim readily follows.
Thus, Theorems 3.10 and 3.12 apply and give completely positive linear maps $\Phi^{\phi}:C^{*}({\mathcal{B}})\to C^{*}({\mathcal{B}})$ and $M^{\phi}:C_{r}^{*}({\mathcal{B}})\to C_{r}^{*}({\mathcal{B}})$ determined by
$\Phi^{\phi}(j_{g}^{{\mathcal{B}}}(b))=\phi(g)j_{g}^{{\mathcal{B}}}(b),\quad M^{\phi}(\lambda_{g}^{{\mathcal{B}}}(b))=\phi(g)\lambda_{g}^{{\mathcal{B}}}(b)$
for all $g\in G$ and $b\in B_{g}$.
Note that if $\phi:G\to{\mathbb{T}}$ is a character, then $T^{\phi}$ is easily seen to be a morphism from ${\mathcal{B}}$ into itself, and the associated map $\Phi^{\gamma}$ (resp. $M^{\gamma}$) is a *-automorphism of $C^{*}({\mathcal{B}})$ (resp. $C_{r}^{*}({\mathcal{B}})$).
Example 4.6.
Let again ${\mathcal{B}}=(B_{g})_{g\in G}$ be Fell bundle. We may then also consider maps of the type used by Exel to introduce the approximation property for Fell bundles [26, 27], cf. Example 5.2. Let $\xi:G\to B_{e}$ be a finitely supported function. For each $g\in G$, define a linear map $T_{g}:B_{g}\to B_{g}$ by
$T_{g}(b)=\sum_{h\in G}\xi(gh)^{*}b\,\xi(h)\quad\text{for all }b\in B_{g}.$
Then $T=(T_{g})_{g\in G}$ is a positive definite ${\mathcal{B}}$-bundle map.
Indeed, let $g_{1},\ldots,g_{n}\in G$ and $b_{i},c_{i}\in B_{g_{i}}$ for $i=1,\ldots,n$. For each $j\in\{1,\ldots,n\}$, let $d_{j}:G\to B_{e}$ be the finitely supported function defined by
$d_{j}(g)=b_{j}\,\xi(g_{j}^{-1}g)\,c_{j}^{*}$
for all $g\in G$, and consider $d_{j}$ as a function in the Hilbert $B_{e}$-module $\ell^{2}(G,B_{e})$. Then for each $i,j\in\{1,\ldots,n\}$ we have
$c_{i}\,T_{g_{i}^{-1}g_{j}}(b_{i}^{*}b_{j})\,c_{j}^{*}=\sum_{h\in G}c_{i}\,\xi(g_{i}^{-1}g_{j}h)^{*}b_{i}^{*}b_{j}\,\xi(h)\,c_{j}^{*}=\sum_{g\in G}d_{i}(g)^{*}d_{j}(g)=\langle d_{i},d_{j}\rangle_{B_{e}}.$
Hence,
$\sum_{i,j=1}^{n}c_{i}\,T_{g_{i}^{-1}g_{j}}(b_{i}^{*}b_{j})\,c_{j}^{*}=\sum_{i,j=1}^{n}\langle d_{i},d_{j}\rangle_{B_{e}}=\big{\langle}\sum_{i=1}^{n}d_{i},\,\sum_{j=1}^{n}d_{j}\big{\rangle}_{B_{e}}\in(B_{e})^{+}.$
Example 4.7.
Let ${\mathcal{A}}=(A_{g})_{g\in G}$ and ${\mathcal{B}}=(B_{h})_{h\in H}$ denote the canonical Fell bundles associated to some discrete $C^{*}$-dynamical systems $\Sigma=(A,G,\alpha)$ and $\Omega=(B,H,\beta)$. For simplicity, we only consider untwisted systems here, although the case where both $\Sigma$ and $\Omega$ are twisted systems can be handled in a similar way. We recall that each fibre $A_{g}:=\{(a,g):a\in A\}$ is identified with $A$ as a Banach space via the map $(a,g)\mapsto a$, the multiplication $A_{g}\times A_{g^{\prime}}\to A_{gg^{\prime}}$ and the involution $A_{g}\to A_{g^{-1}}$ being given by
$(a,g)(a^{\prime},g^{\prime}):=(a\alpha_{g}(a^{\prime}),gg^{\prime}),$
and
$(a,g)^{*}:=(\alpha_{g^{-1}}(a^{*}),g^{-1}),$
respectively. Each $B_{h}=\{(b,h):b\in B\}$ is defined similarly, along with the Fell bundle operations.
We also recall that $C^{*}({\mathcal{A}})$ (resp. $C_{r}^{*}({\mathcal{A}})$) corresponds to the full (resp. reduced) $C^{*}$-crossed product $C^{*}(\Sigma)$ (resp. $C_{r}^{*}(\Sigma)$) associated to $\Sigma$, and similarly with $C^{*}({\mathcal{B}})$ (resp. $C_{r}^{*}({\mathcal{B}})$). For $g\in G$ and $a\in A$, we will identify $\widehat{j}_{g}^{\mathcal{A}}((a,g))$ with the canonical element $a\,U_{\Sigma}(g)$ in $C^{*}(\Sigma)$, and $\lambda_{g}^{\mathcal{A}}((a,g))$ with the canonical element $a\,\lambda_{\Sigma}(g)$ in $C_{r}^{*}(\Sigma)$.
Consider a family $(\widetilde{T}_{g})_{g\in G}$ of bounded linear maps from $A$ to $B$. For each $g\in G$, define a bounded linear map $T_{g}:A_{g}\to B_{\varphi(g)}$ by
$T_{g}((a,g))=\big{(}\widetilde{T}_{g}(a),\varphi(g))$
for all $a\in A$. Then $T=(T_{g})_{g\in G}$ is easily seen to be an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. Moreover, $T$ is positive definite if and only if the family $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-$\varphi$-$\Omega$ positive definite in the following sense: for every $g_{1},\ldots,g_{n}\in G$ and $a_{1},\ldots,a_{n}\in A$, the matrix
$\Big{[}\beta_{\varphi(g_{i})}\big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}(\alpha_{{g_{i}}^{-1}}(a_{i}^{*}a_{j}))\big{)}\Big{]}$
is positive in $M_{n}(B)$. [^1] When $G=H$, $\varphi={\rm id}_{G}$, $\Sigma=\Omega$ and $A$ is unital, this amounts to say that $(\tilde{T}_{g})_{g\in G}$ is positive definite w.r.t. $\Sigma$ in the sense of [9].
Indeed, by definition, $T=(T_{g})_{g\in G}$ is positive definite if and only if
$\sum_{i,j=1}^{n}(b_{i},\varphi(g_{i}))\,T_{g_{i}^{-1}g_{j}}\big{(}(a_{i},g_{i})^{*}(a_{j},g_{j})\big{)}\,(b_{j},\varphi(g_{j}))^{*}\,\in\,B^{+}\times\{e\}$
whenever $g_{1},\ldots,g_{n}\in G$, $a_{1},\ldots,a_{n}\in A$ and $b_{1},\ldots,b_{n}\in B$. On the other hand, $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-$\varphi$-$\Omega$ positive definite if and only if
$\sum_{i,j=1}^{n}b_{i}\,\beta_{\varphi(g_{i})}\big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}\big{(}\alpha_{{g_{i}}^{-1}}({a_{i}}^{*}a_{j})\big{)}\big{)}\,{b_{j}}^{*}\,\in B^{+}$
whenever $g_{1},\ldots,g_{n}\in G$ and $a_{1},\ldots,a_{n}\in A$ and $b_{1},\ldots,b_{n}\in B$. As we have
$~(b_{i},\varphi(g_{i}))$ $\,T_{g_{i}^{-1}g_{j}}\big{(}(a_{i},g_{i})^{*}(a_{j},g_{j})\big{)}\,(b_{j},\varphi(g_{j}))^{*}$
$=(b_{i},\varphi(g_{i}))\,T_{g_{i}^{-1}g_{j}}\big{(}(\alpha_{{g_{i}}^{-1}}({a_{i}}^{*}),{g_{i}}^{-1})(a_{j},g_{j})\big{)}\,(\beta_{{\varphi(g_{j})}^{-1}}({b_{j}}^{*}),\varphi({g_{j}})^{-1})$
$=(b_{i},\varphi(g_{i}))\,T_{g_{i}^{-1}g_{j}}\big{(}(\alpha_{{g_{i}}^{-1}}({a_{i}}^{*})\alpha_{{g_{i}}^{-1}}(a_{j}),{g_{i}}^{-1}g_{j})\big{)}\,(\beta_{{\varphi(g_{j})}^{-1}}({b_{j}}^{*}),\varphi({g_{j}})^{-1})$
$=(b_{i},\varphi(g_{i}))\,T_{g_{i}^{-1}g_{j}}\big{(}(\alpha_{{g_{i}}^{-1}}({a_{i}}^{*}a_{j}),{g_{i}}^{-1}g_{j})\big{)}\,(\beta_{{\varphi(g_{j})}^{-1}}({b_{j}}^{*}),\varphi({g_{j}})^{-1})$
$=(b_{i},\varphi(g_{i}))\,\big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}\big{(}\alpha_{{g_{i}}^{-1}}({a_{i}}^{*}a_{j})\big{)},\varphi({g_{i}}^{-1}g_{j})\big{)}\,(\beta_{{\varphi(g_{j})}^{-1}}({b_{j}}^{*}),\varphi({g_{j}})^{-1})$
$=\big{(}b_{i}\,\beta_{\varphi(g_{i})}\big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}\big{(}\alpha_{{g_{i}}^{-1}}({a_{i}}^{*}a_{j})\big{)}\big{)},\varphi(g_{j})\big{)}\,(\beta_{{\varphi(g_{j})}^{-1}}({b_{j}}^{*}),\varphi({g_{j}})^{-1})$
$=\big{(}b_{i}\,\beta_{\varphi(g_{i})}\big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}\big{(}\alpha_{{g_{i}}^{-1}}({a_{i}}^{*}a_{j})\big{)}\big{)}\,{b_{j}}^{*},e\big{)},$
the claim readily follows.
Applying Theorem 3.10, we get that $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-$\varphi$-$\Omega$ positive definite if and only if there exists a completely positive map $\Phi_{\widetilde{T}}:C^{*}(\Sigma)\to C^{*}(\Omega)$ such that
$\Phi_{\widetilde{T}}(aU_{\Sigma}(g))=\widetilde{T}_{g}(a)U_{\Omega}(\varphi(g))$
for all $a\in A$ and $g\in G$, in which case we have $\|\Phi_{\widetilde{T}}\|=\|\widetilde{T}_{e}\|$.
Moreover, if we assume that ${\rm ker}(\varphi)$ is amenable, then Theorem 3.12 gives that $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-$\varphi$-$\Omega$ positive definite if and only if there exists a completely positive map $M_{\widetilde{T}}:C_{r}^{*}(\Sigma)\to C_{r}^{*}(\Omega)$ such that
$M_{\widetilde{T}}(a\lambda_{\Sigma}(g))=\widetilde{T}_{g}(a)\lambda_{\Omega}(\varphi(g))$
for all $a\in A$ and $g\in G$, in which case $\|M_{\widetilde{T}}\|=\|\widetilde{T}_{e}\|$.
We note that in the case where $G=H$, $\varphi={\rm id}_{G}$, $\Sigma=\Omega$ and $A$ is unital, these two results boil down to the equivalence of (b), (c) and (d) in [9, Corollary 4.4]. Under the same assumptions, we also illustrated in [9, Example 4.1] how positive definite families $(\widetilde{T}_{g})_{g\in G}$ may be constructed from equivariant representations of $\Sigma$. We will now describe how such families arise when $\Omega$ is possibly different from $\Sigma$ and $A$ may be non-unital, still assuming that $G=H$ and $\varphi={\rm id}_{G}$.
Let $X$ be a right Hilbert $B$-module such that $A$ acts on $X$ from the left by adjointable operators, and assume $\gamma$ is a homomorphism from $G$ into the group of ${\mathbb{C}}$-linear invertible isometries of $X$ satisfying
  • (i)
    $\gamma_{g}(a\cdot x)=\alpha_{g}(a)\cdot\gamma_{g}(x)$
  • (ii)
    $\gamma_{g}(x\cdot b)=\gamma_{g}(x)\cdot\beta_{g}(b)$
  • (iii)
    $\langle\gamma_{g}(x),\gamma_{g}(y)\rangle_{B}=\beta_{g}(\langle x,y\rangle_{B})$
for each $g\in G,a\in A,b\in B$ and $x,y\in X$. (This means that $\gamma$ is an $\alpha$-$\beta$ compatible action of $G$ on the right Hilbert $A$-$B$-bimodule $X$ in the sense of [24], except that we don’t assume non-degeneracy of the left action of $A$ on $X$.)
For each $g\in G$ define a linear map $\widetilde{T}_{g}:A\to B$ by
$\widetilde{T}_{g}(a)=\big{\langle}x,a\cdot(\gamma_{g}(x))\big{\rangle}_{B}$
for all $a\in A$. Then $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-${\rm id}_{G}$-$\Omega$ positive-definite. Indeed, let $g_{1},\ldots,g_{n}\in G$ and $a_{1},\ldots,a_{n}\in A$. Then for $i,j\in\{1,\ldots,n\}$ we have
$\beta_{g_{i}}\Big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}\big{(}\alpha_{{g_{i}}^{-1}}(a_{i}^{*}a_{j})\big{)}\Big{)}$ $=\beta_{g_{i}}\Big{(}\big{\langle}x,\alpha_{{g_{i}}^{-1}}(a_{i}^{*}a_{j})\cdot(\gamma_{g_{i}^{-1}g_{j}}(x))\big{\rangle}_{B}\Big{)}$
$=\big{\langle}\gamma_{g_{i}}(x),\gamma_{g_{i}}\big{(}\alpha_{{g_{i}}^{-1}}(a_{i}^{*}a_{j})\cdot\gamma_{g_{i}^{-1}g_{j}}(x)\big{)}\big{\rangle}_{B}$
$=\big{\langle}\gamma_{g_{i}}(x),\alpha_{g_{i}}(\alpha_{{g_{i}}^{-1}}(a_{i}^{*}a_{j}))\cdot\gamma_{g_{i}}(\gamma_{g_{i}^{-1}g_{j}}(x))\big{\rangle}_{B}$
$=\big{\langle}\gamma_{g_{i}}(x),(a_{i}^{*}a_{j})\cdot\gamma_{g_{j}}(x)\big{\rangle}_{B}$
$=\big{\langle}a_{i}\cdot\gamma_{g_{i}}(x),a_{j}\cdot\gamma_{g_{j}}(x)\big{\rangle}_{B}$
$=\big{\langle}y_{i},y_{j}\rangle_{B}$
where $y_{i}:=a_{i}\cdot\gamma_{g_{i}}(x)\in X$ for every $i$. Thus,
$\Big{[}\beta_{g_{i}}\big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}(\alpha_{{g_{i}}^{-1}}(a_{i}^{*}a_{j}))\big{)}\Big{]}=\big{[}\big{\langle}y_{i},y_{j}\rangle_{B}\big{]}\,\in M_{n}(B)^{+},$
as desired.
We note that by adapting the proof of [9, Theorem 4.5], one can show that that every $\Sigma$-${\rm id}_{G}$-$\Omega$ positive definite family may be obtained in such a way, at least when $A$ and $B$ are unital.
Example 4.8.
Assume ${\mathcal{B}}=(B_{g})_{g\in G}$ is a Fell sub-bundle of a Fell Bundle ${\mathcal{A}}=(A_{g})_{g\in G}$, as defined in [27, Definition 21.5], and $E=(E_{g})$ is a conditional expectation from ${\mathcal{A}}$ to ${\mathcal{B}}$ in the sense of Exel [27, Definition 21.19]. Thus, each $E_{g}$ is a bounded idempotent linear mapping from $A_{g}$ onto $B_{g}$, and we have
  • (i)
    $E_{e}$ is a conditional expectation from $A_{e}$ onto $B_{e}$,
  • (ii)
    for each $g\in G$ and $a\in A_{g}$, we have $E_{g}(a)^{*}=E_{g^{-1}}(a^{*})$, and
  • (iii)
    for each $g,h\in G$ and $a\in A_{g},b\in B_{h}$, we have
    $E_{gh}(ab)=E_{g}(a)\,b,\quad E_{hg}(ba)=bE_{g}(a).$
Then $E$ is a positive definite ${\mathcal{A}}$$-{\mathcal{B}}$ bundle map.
Indeed, if $g_{1},\ldots,g_{n}\in G$, $a_{i}\in A_{g_{i}}$ and $b_{i}\in B_{g_{i}}$ for $i=1,\ldots,n$, and we set $c:=\sum_{j=1}^{n}a_{j}b_{j}^{*}\in A_{e}$, then
$\sum_{i,j=1}^{n}b_{i}\,E_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\,b_{j}^{*}$ $=\sum_{i,j=1}^{n}\,E_{g_{i}g_{i}^{-1}g_{j}g_{j}^{-1}}(b_{i}a_{i}^{*}a_{j}b_{j}^{*})$
$=\sum_{i,j=1}^{n}\,E_{e}(b_{i}a_{i}^{*}a_{j}b_{j}^{*})=E_{e}(c^{*}c)$
which is positive in $B_{e}$ since $E_{e}$ is a positive map.
Applying again Theorem 3.10 and Theorem 3.12, we obtain the existence of completely positive linear maps $\Phi_{E}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ and $M_{E}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ satisfying that
$\Phi_{E}(\widehat{j}^{\,{\mathcal{A}}}_{g}(a))=\widehat{j}^{\,{\mathcal{B}}}_{g}(E_{g}(a)),\quad M_{E}(\lambda^{\mathcal{A}}_{g}(a))=\lambda^{\mathcal{B}}_{g}(E_{g}(a))$
for all $g\in G$ and $a\in A_{g}$. It is well-known that such a map $M_{E}$ exists and is a conditional expectation when $C_{r}^{*}({\mathcal{B}})$ is identified with its canonical copy inside $C_{r}^{*}({\mathcal{A}})$ (cf. [27, Theorem 21.29]).
Further, since ${\mathcal{B}}$ admits a conditional expectation in the sense of Exel (by assumption), we can identify $C^{*}({\mathcal{B}})$ with its canonical copy inside $C^{*}({\mathcal{A}})$ (cf. [27, Theorem 21.30]). Then, for every $g\in G$ and $b\in B_{g}$, we get that
$\Phi_{E}(\widehat{j}^{\,{\mathcal{A}}}_{g}(b))=\widehat{j}^{\,{\mathcal{A}}}_{g}(E_{g}(b))=\widehat{j}^{\,{\mathcal{A}}}_{g}(b).$
Hence, by continuity of $\Phi_{E}$, we get that $\Phi_{E}(x)=x$ for every $x\in C^{*}({\mathcal{B}})$, i.e., $\Phi_{E}$ is a projection from $C^{*}({\mathcal{A}})$ onto $C^{*}({\mathcal{B}})$. Since $\|\Phi_{E}\|=\|E_{e}\|=1$, we get that $\Phi_{E}$ is a conditional expectation.
Example 4.9.
Let $Y$ be a right Hilbert $C_{r}^{*}({\mathcal{B}})$-module and $y\in Y$. Assume that $C_{r}^{*}({\mathcal{A}})$ acts on $Y$ from the left by adjointable operators. For each $g\in G$, define a linear map $T_{g}:A_{g}\to B_{\varphi(g)}$ by
$T_{g}(a)=E^{\mathcal{B}}_{\varphi(g)}\Big{(}\big{\langle}y,\lambda^{\mathcal{A}}_{g}(a)y\big{\rangle}_{C_{r}^{*}({\mathcal{B}})}\Big{)}\quad\text{for all }a\in A_{g}.$ (5)
Then $T=(T_{g})$ is clearly an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. Moreover, it is positive definite:
Indeed, let $g_{1},\ldots,g_{n}\in G$ and $a_{i}\in A_{g_{i}},b_{i}\in B_{\varphi(g_{i})}$ for $i=1,\ldots,n$. Then we have
$b_{i}\,T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\,b_{j}^{*}$ $=b_{i}\,E^{\mathcal{B}}_{\varphi(g_{i}^{-1}g_{j})}\big{(}\langle y,\lambda^{\mathcal{A}}_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})y\rangle_{C_{r}^{*}({\mathcal{B}})}\big{)}\,b_{j}^{*}$
$=b_{i}\,E^{\mathcal{B}}_{\varphi(g_{i})^{-1}\varphi(g_{j})}\big{(}\langle\lambda^{\mathcal{A}}_{g_{i}}(a_{i})y,\lambda^{\mathcal{A}}_{g_{j}}(a_{j})y\rangle_{C_{r}^{*}({\mathcal{B}})}\big{)}b_{j}^{*}$
$=b_{i}\,E^{\mathcal{B}}_{\varphi(g_{i})^{-1}\varphi(g_{j})}\big{(}z_{ij}\big{)}b_{j}^{*},$
where $z_{ij}:=\langle\lambda^{\mathcal{A}}_{g_{i}}(a_{i})y,\lambda^{\mathcal{A}}_{g_{j}}(a_{j})y\rangle_{C_{r}^{*}({\mathcal{B}})}$ for each $i,j$. Now, using [27, Proposition 17.12], we get that
$b_{i}\,E^{\mathcal{B}}_{\varphi(g_{i})^{-1}\varphi(g_{j})}\big{(}z_{ij}\big{)}b_{j}^{*}=\lambda^{\mathcal{B}}_{g_{i}}(b_{i})\,z_{ij}\,\lambda^{\mathcal{B}}_{g_{j}}(b_{j})^{*}$
for each $i,j$. Since $Z:=[z_{ij}]\in M_{n}(C_{r}^{*}({\mathcal{B}}))^{+}$, we get that
$\sum_{i,j=1}^{n}b_{i}\,T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\,b_{j}^{*}=\sum_{i,j=1}^{n}\lambda^{\mathcal{B}}_{g_{i}}(b_{i})\,z_{ij}\,\lambda^{\mathcal{B}}_{g_{j}}(b_{j})^{*}=\big{\langle}\Omega,Z\,\Omega\big{\rangle}\in C_{r}^{*}({\mathcal{B}})^{+}\,,$
[…]
We leave to the reader to check that one obtains a similar result if one replaces the reduced cross-sectional $C^{*}$-algebras by their full counterparts and defines $T_{g}:A_{g}\to B_{\varphi(g)}$ by
$T_{g}(a)=\big{(}E^{\mathcal{B}}_{\varphi(g)}\circ\Lambda^{\mathcal{B}}\big{)}\Big{(}\big{\langle}y,\widehat{j}^{\mathcal{A}}_{g}(a)y\big{\rangle}_{C^{*}({\mathcal{B}})}\Big{)}\quad\text{for all }a\in A_{g}.$ (6)
Conversely, let us start with a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$. We can then form the completely positive map $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$. Following Paschke [55, Section 5] (see also [46, p. 48]), we may then construct from $\Phi_{T}$ a right Hilbert $C^{*}({\mathcal{B}})$-module $Y$ and a left action of $C^{*}({\mathcal{A}})$ on $Y$ by adjointable operators. Assume that $A_{e}$ and $B_{e}$ are unital. Then $C^{*}({\mathcal{A}})$ and $C^{*}({\mathcal{B}})$ are unital, so [55, Theorem 5.2] gives that there exists some $y\in Y$ such that
$\Phi_{T}(X)=\langle y,\,X\cdot y\rangle_{C^{*}({\mathcal{B}})}$
for all $X\in C^{*}({\mathcal{A}})$. Then for all $g\in G$ and $a\in A_{g}$ we have
$\langle y,\,\widehat{j}^{\,{\mathcal{A}}}_{g}(a)\cdot y\rangle_{C^{*}({\mathcal{B}})}=\Phi_{T}(\widehat{j}^{\,{\mathcal{A}}}_{g}(a))=\widehat{j}_{\varphi(g)}^{\,{\mathcal{B}}}(T_{g}(a)),$
hence
$\big{(}E^{\mathcal{B}}_{\varphi(g)}\circ\Lambda^{\mathcal{B}}\big{)}\big{(}\langle y,\widehat{j}^{\,{\mathcal{A}}}_{g}(a)\cdot y\rangle_{C^{*}({\mathcal{B}})}\big{)}=E^{\mathcal{B}}_{\varphi(g)}\big{(}\Lambda^{\mathcal{B}}\big{(}\widehat{j}_{\varphi(g)}^{\,{\mathcal{B}}}(T_{g}(a))\big{)}\big{)}=E^{\mathcal{B}}_{\varphi(g)}\big{(}\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\big{)}=T_{g}(a).$
This shows that every positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T$ may be written in the form (6). If we further assume that ${\rm ker}(\varphi)$ is amenable, then arguing analogously with the completely positive map $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$, we also get that every such bundle map $T$ is of the form (5).
5 An approximation property for Fell bundles
Recall that a discrete group $G$ is amenable if and only if there exists a net $\{\varphi_{i}\}$ of normalized finitely supported positive definite functions on $G$ such that $\varphi_{i}\to 1$ pointwise (see e.g. [14, Theorem 2.6.8]). An analogous property for Fell bundles is as follows.
Definition 5.1.
A Fell Bundle ${\mathcal{B}}=(B_{g})_{g\in G}$ over a discrete group $G$ has the PD-approximation property if there exists a net $\{T^{i}\}_{i\in I}$ of ${\mathcal{B}}$-bundle maps satisfying the following properties:
  • (i)
    For each $i\in I$, $T^{i}=(T^{i}_{g})_{g\in G}$ is positive definite and its support $\{g\in G:T^{i}_{g}\neq 0\}$ is finite.
  • (ii)
    The net $\{T^{i}\}_{i\in I}$ is uniformly bounded in the sense that $\sup_{i}\|T^{i}_{e}\|<\infty$.
  • (iii)
    $\lim_{i}\|T^{i}_{g}(b)-b\|=0$ for every $g\in G$ and $b\in B_{g}$.
Example 5.2.
Recall from [26, 27] that a Fell bundle ${\mathcal{B}}=(B_{g})_{g\in G}$ is said to have the Exel approximation property (AP) whenever there exists a net $\{\xi_{i}\}_{i\in I}$ of finitely supported functions from $G$ to $B_{e}$ satisfying that
  • $\sup_{i\in I}\big{\|}~\sum_{h\in G}\xi_{i}(h)^{*}\xi_{i}(h)\big{\|}<\infty$,
  • $\lim_{i}\|\sum_{h\in G}\xi_{i}(gh)^{*}b\,\xi_{i}(h)-b\|=0$ for every $g\in G$ and $b\in B_{g}$.
It has recently been shown that the AP is equivalent to several other properties, such as AD-amenability, see [3, 16, 18]. It can be also extended to Fell bundles over groupoids, cf. [38].
We note that a Fell bundle ${\mathcal{B}}=(B_{g})_{g\in G}$ has the PD-approximation property whenever ${\mathcal{B}}$ has the AP. Indeed, assume that there exists a net $\{\xi_{i}\}_{i\in I}$ as above. For each $i\in I$ and $g\in G$, let then $T^{i}_{g}:B_{g}\to B_{g}$ be the map defined by
$T^{i}_{g}(b)=\sum_{h\in G}\xi_{i}(gh)^{*}b\,\xi_{i}(h)\quad\text{for all }b\in B_{g}$
As explained in Example 4.6, each $T^{i}=(T^{i}_{g})_{g\in G}$ is a positive definite ${\mathcal{B}}$-bundle map, and one readily sees that $\{T^{i}\}_{i\in I}$ is a net guaranteeing that ${\mathcal{B}}$ has the PD-approximation property.
This implies in particular that ${\mathcal{B}}=(B_{g})_{g\in G}$ has the PD-approximation property whenever $G$ is amenable.
Example 5.3.
Let $\Sigma=(A,G,\alpha)$ be a (unital) discrete $C^{*}$-dynamical system, and let ${\mathcal{B}}$ denote the associated Fell bundle over $G$. If the system $\Sigma$ is amenable in the sense of [10], then, making use of Example 4.7, we get that ${\mathcal{B}}$ has the PD-approximation property.
Recall that a Fell bundle ${\mathcal{B}}$ is said to have the weak containment property (WCP), or to be Exel-amenable, whenever the canonical map $\Lambda^{\mathcal{B}}:C^{*}({\mathcal{B}})\to C_{\rm r}^{*}({\mathcal{B}})$ is injective.
Theorem 5.4.
Assume ${\mathcal{B}}=(B_{g})_{g\in G}$ has the PD-approximation property. Then ${\mathcal{B}}$ has the WCP. Hence $C^{*}({\mathcal{B}})\simeq C_{r}^{*}({\mathcal{B}})$.
Proof.
The proof of the first assertion is an adaptation of the proof for the case of group $C^{*}$-algebras associated to amenable groups given in [14, Theorem 2.6.8] (see also the proof of [10, Theorem 4.6]).
We start by observing that if $F$ is a finite subset of $G$ and $C_{F}$ is the subspace of $C_{c}({\mathcal{B}})$ given by $C_{F}=\{f\in C_{c}({\mathcal{B}}):{\rm supp}(f)\subseteq F\}$, then we have that $\kappa(C_{F})$ is a closed subspace of $C^{*}({\mathcal{B}})$. Here we write $\kappa$ for the canonical map $\kappa^{\mathcal{B}}:C_{c}({\mathcal{B}})\to C^{*}({\mathcal{B}})$.
Indeed, let $\{f_{n}\}$ be a sequence in $C_{F}$ such that $\{\kappa(f_{n})\}$ converges to some $x\in C^{*}({\mathcal{B}})$. For each $g\in G$, let $\widetilde{E}_{g}:C^{*}({\mathcal{B}})\to B_{g}$ be the linear map defined by $\widetilde{E}_{g}:=E^{\mathcal{B}}_{g}\circ\Lambda^{\mathcal{B}}$, which is clearly continuous. For each $g\in F$, we then have
$f_{n}(g)=\widetilde{E}_{g}(\kappa(f_{n}))\to\widetilde{E}_{g}(x)\in B_{g}\,\,\text{as}\,\,n\to\infty\,.$
Thus, letting $f\in C_{F}$ be defined by $f=\sum_{g\in F}\widetilde{E}_{g}(x)\odot g$, and using (1), we get
$\|x-\kappa(f)\|_{u}$ $\leq\|x-\kappa(f_{n})\|_{u}+\|\kappa(f_{n}-f)\|_{u}$
$\leq\|x-\kappa(f_{n})\|_{u}+\sum_{g\in F}\|f_{n}(g)-\widetilde{E}_{g}(x)\|\to 0\,\,\text{as}\,\,n\to\infty\,.$
Hence, $x=\kappa(f)\in\kappa(C_{F})$, as desired.
Next, let $x\in C^{*}({\mathcal{B}})$. Consider a positive definite ${\mathcal{B}}$-bundle map $T$ having finite support. Then $\Phi_{T}(x)\in\kappa(C_{c}({\mathcal{B}}))$: indeed, letting $F$ denote the support of $T$ and $\{f_{n}\}$ be a sequence in $C_{c}({\mathcal{B}})$ such that $\{\kappa(f_{n})\}$ converges to $x$, we have
$\Phi_{T}(\kappa(f_{n}))=\kappa(\phi_{T}(f_{n}))=\kappa\Big{(}\sum_{g\in F}\big{(}T_{g}\big{(}f_{n}(g)\big{)}\odot g\Big{)}\,\in\kappa(C_{F})$
for each $n$, so $\Phi_{T}(x)=\lim_{n}\Phi_{T}(\kappa(f_{n}))$ belongs to $\overline{\kappa(C_{F})}=\kappa(C_{F})\subseteq\kappa(C_{c}({\mathcal{B}}))$.
Assume now that $\Lambda(x)=0$. Let $\{T^{i}\}_{i\in I}$ be a net as guaranteed by the PD-approximation property of ${\mathcal{B}}$. Theorem 3.10 (resp. Theorem 3.12) gives a net $\{\Phi_{T^{i}}\}_{i\in I}$ (resp. $\{M_{T^{i}}\}_{i\in I}$) of completely positive maps on $C^{*}({\mathcal{B}})$ (resp.  $C_{r}^{*}({\mathcal{B}})$). Note that for each $i\in I$, we have $\Lambda\circ\Phi_{T^{i}}=M_{T^{i}}\circ\Lambda$ on $C^{*}({\mathcal{B}})$, since the maps on both sides are continuous and agree on $\kappa(C_{c}({\mathcal{B}}))$. Hence, for each $i\in I$, we get
$\Lambda\big{(}\Phi_{T^{i}}(x)\big{)}=M_{T^{i}}\big{(}\Lambda(x)\big{)}=0\,.$
Since each $T^{i}$ is finitely supported, we have $\Phi_{T^{i}}(x)\in\kappa(C_{c}({\mathcal{B}}))$, as established above. But $\Lambda$ is injective on $\kappa(C_{c}({\mathcal{B}}))$, cf. [27, Proposition 17.9], so we obtain that $\Phi_{T^{i}}(x)=0$ for each $i\in I$.
In order to conclude that $x=0$, we will use the observation that for every $y\in C^{*}({\mathcal{B}})$ we have $y=\lim_{i}\Phi_{T^{i}}(y)$. To see this, let $f\in C_{c}({\mathcal{B}})$ have support $F\subseteq G$. Using (1), we get that
$\|\Phi_{T^{i}}(\kappa(f))-\kappa(f)\|_{\rm u}=\Big{\|}\sum_{g\in F}\,\kappa\Big{(}\big{[}T_{g}^{i}(f(g))-f(g)\big{]}\odot g\Big{)}\Big{\|}_{\rm u}\,\leq\,\sum_{g\in F}\,\big{\|}T_{g}^{i}(f(g))-f(g)\|\,.$
Hence, using the assumption that $\lim_{i}\|T^{i}_{g}(b)-b\|=0$ for every $g\in G$ and $b\in B_{g}$, we get that $\lim_{i}\|\Phi_{T^{i}}(\kappa(f))-\kappa(f)\|_{\rm u}=0$. This implies that $\lim_{i}\|\Phi_{T^{i}}(y)-y\|_{\rm u}=0$ for all $y\in\kappa\big{(}C_{c}({\mathcal{B}})\big{)}$. If $\{u_{j}\}_{j\in J}$ is an approximate unit for $A$, then $\{\kappa\big{(}u_{j}\odot e)\}_{j\in J}$ is easily seen to be an approximate unit for $C^{*}({\mathcal{B}})$, so we get that
$\|\Phi_{T^{i}}\|=\lim_{j}\|\Phi_{T^{i}}\big{(}\kappa(u_{j}\odot e)\big{)}\|_{\rm u}=\lim_{j}\|\kappa\big{(}T_{e}^{i}(u_{j})\odot e\big{)}\|_{\rm u}=\lim_{j}\|T_{e}^{i}(u_{j})\|=\|T_{e}^{i}\|$
for each $i\in I$. By property (ii) for $\{T^{i}\}_{i\in I}$, we have $\sup_{i\in I}\|T^{i}_{e}\|<\infty$, so we get that $\sup_{i\in I}\|\Phi_{T^{i}}\|<\infty$. Hence, since $\kappa\big{(}C_{c}({\mathcal{B}})\big{)}$ is dense in $C^{*}({\mathcal{B}})$, an $\varepsilon/3$-argument gives that $\lim_{i}\Phi_{T^{i}}(y)=y\,$ for every $y\in C^{*}({\mathcal{B}})$.
Applying this observation to $x$, we get that $x=\lim_{i}\Phi_{T^{i}}(x)=0$. This shows that $\Lambda$ is injective, as desired. $\square$
Our next result involves the tensor product Fell bundle $C\otimes_{\rm min}{\mathcal{B}}=\big{(}C\otimes_{\rm min}B_{g}\big{)}_{g\in G}$, which is defined for any $C^{*}$-algebra $C$ in [27, Definition 25.4]. A similar result when ${\mathcal{B}}$ has the AP is shown by Exel in [27, Proposition 25.9].
Proposition 5.5.
Assume ${\mathcal{B}}=(B_{g})_{g\in G}$ is a Fell bundle having the PD-approximation property, and let $C$ be a $C^{*}$-algebra. Then $C\otimes_{\rm min}{\mathcal{B}}$ has the PD-approximation property.
Proof.
We recall from [27, Sections 24 and 25] that the Fell bundle $\mathcal{D}:=C\otimes_{\rm min}{\mathcal{B}}$ is the completion of the pre-Fell-bundle obtained by equipping the algebraic Fell bundle $C\odot{\mathcal{B}}=(C\odot B_{g})_{g\in G}$ with the minimal $C^{*}$-norm $\|\cdot\|_{\rm min}$ on $C\odot B_{e}$. The norm on each fiber $C\odot B_{g}$ is then given by $\|d\|=\|d^{*}d\|_{\rm min}^{1/2}$ for $d\in C\odot B_{g}$, and the fiber of $C\otimes_{\rm min}{\mathcal{B}}$ at $g$, denoted by $C\otimes_{\rm min}B_{g}$, is the completion of $C\odot B_{g}$ w.r.t. this norm.
Let $\{T^{i}\}_{i\in I}$ be a net as guaranteed by the PD-approximation property of ${\mathcal{B}}$ and consider $i\in I$. Then by Theorem 3.12, $T^{i}$ is a reduced ${\mathcal{B}}$-bundle map such that $M_{T^{i}}$ is completely positive on $C_{r}^{*}({\mathcal{B}})$. Letting ${\rm id}_{C}$ denote the identity map on $C$ and using [14, Theorem 3.5.3], we get a completely positive map $\widetilde{M}_{i}:={\rm id}_{C}\otimes M_{T^{i}}$ on $C\otimes_{\rm min}C_{r}^{*}({\mathcal{B}})$ satisfying
$\widetilde{M}_{i}(c\otimes\lambda^{\mathcal{B}}_{g}(b))=c\otimes M_{T^{i}}(\lambda^{\mathcal{B}}_{g}(b))=c\otimes\lambda^{\mathcal{B}}_{g}(T^{i}_{g}(b))$
for all $c\in C,g\in G$ and $b\in B_{g}$, and $\|\widetilde{M}_{i}\|=\|{\rm id}_{C}\|~\|M_{T^{i}}\|=\|T^{i}_{e}\|$. Now, using [27, Theorem 25.8], there is a *-isomorphism $\psi:C_{r}^{*}(\mathcal{D})\to C\otimes_{\rm min}C_{r}^{*}({\mathcal{B}})$ satisfying
$\psi(\lambda^{\mathcal{D}}_{g}(c\otimes b))=c\otimes\lambda^{\mathcal{B}}_{g}(b)$
for all $c\in C,g\in G$ and $b\in B_{g}$.
We therefore get a completely positive map $\Psi_{i}:=\psi^{-1}\circ\widetilde{M}_{i}\circ\psi$ on $C_{r}^{*}(\mathcal{D})$ satisfying that
$\Psi_{i}(\lambda^{\mathcal{D}}_{g}(c\otimes b))=\lambda^{\mathcal{D}}_{g}(c\otimes T^{i}_{g}(b))$
for all $c\in C,g\in G$ and $b\in B_{g}$, and $\|\Psi_{i}\|=\|T^{i}_{e}\|$. Hence, using Proposition 3.7, we get that $S^{i}:=(S^{i}_{g})_{g\in G}$, where for each $g\in G$, $S^{i}_{g}:C\otimes_{\rm min}B_{g}\to C\otimes_{\rm min}B_{g}$ is defined by
$S^{i}_{g}:=E^{\mathcal{D}}_{g}\circ\Psi_{i}\circ\lambda^{\mathcal{D}}_{g},$
is a positive definite $\mathcal{D}$-bundle map.
It is now easy to check that the net $\{S^{i}\}_{i\in I}$ satisfies properties (i) and (ii) needed for the PD-approximation property of $\mathcal{D}=C\otimes_{\rm min}{\mathcal{B}}$. To verify that (iii) also holds, using that this net is uniformly bounded, we only have to check that for $c\in C,g\in G$ and $b\in B_{g}$, we have
$\lim_{i}\|S^{i}_{g}(c\otimes b)-c\otimes b\|~=0.$
Now, as $S^{i}_{g}(c\otimes b)=E^{\mathcal{D}}_{g}(\Psi_{i}(\lambda^{\mathcal{D}}_{g}(c\otimes b))=E^{\mathcal{D}}_{g}(\lambda^{\mathcal{D}}_{g}(c\otimes T^{i}_{g}(b)))=c\otimes T^{i}_{g}(b)$, we get that
$\|S^{i}_{g}(c\otimes b)-c\otimes b\|$ $=\|(c\otimes(T^{i}_{g}(b)-b))^{*}(c\otimes(T^{i}_{g}(b)-b))\|_{\rm min}^{1/2}$
$=\|c^{*}c\,\otimes(T^{i}_{g}(b)-b)^{*}(T^{i}_{g}(b)-b)\|_{\rm min}^{1/2}$
$=\|c^{*}c\|^{1/2}\,\|(T^{i}_{g}(b)-b)^{*}(T^{i}_{g}(b)-b)\|^{1/2}$
$=\|c\|\,\|T^{i}_{g}(b)-b\|\to_{i}0,$
as desired. $\square$
Theorem 5.6.
Assume that ${\mathcal{B}}=(B_{g})_{g\in G}$ is a Fell bundle having the $PD$-approximation property. Then $C^{*}({\mathcal{B}})\simeq C_{\rm r}^{*}({\mathcal{B}})$ is nuclear if and only if $B_{e}$ is nuclear.
Proof.
By Theorem 5.4, we know that $C^{*}({\mathcal{B}})\simeq C_{\rm r}^{*}({\mathcal{B}})$. Assume that $B_{e}$ is nuclear. In [27, Proposition 25.10], Exel shows that $C^{*}({\mathcal{B}})\simeq C_{\rm r}^{*}({\mathcal{B}})$ is then nuclear whenever ${\mathcal{B}}$ has the AP. It is straightforward to see that his proof goes through verbatim if one just invokes Proposition 5.5 instead of [27, Proposition 25.9]. Conversely, nuclearity of $B_{e}$ is necessary for the nuclearity of $C_{r}^{*}({\mathcal{B}})$ since there exists a conditional expectation from $C_{r}^{*}({\mathcal{B}})$ onto $B_{e}$. $\square$
We can now include the PD-approximation property in [3, Proposition 7.2] (see also [5, Théorème 4.5], [27, Proposition 25.10] and [16, Corollary 4.23]).
Corollary 5.7.
Let ${\mathcal{B}}=(B_{g})_{g\in G}$ be a Fell bundle. Then the following conditions are equivalent:
  • (i)
    ${\mathcal{B}}$ has the PD-approximation property and $B_{e}$ is nuclear.
  • (ii)
    ${\mathcal{B}}$ has the AP and $B_{e}$ is nuclear.
  • (iii)
    $C^{*}({\mathcal{B}})$ is nuclear.
  • (iv)
    $C_{r}^{*}({\mathcal{B}})$ is nuclear.
Proof.
For the equivalence between ($ii$), ($iii$) and ($iv$), see [3, Proposition 7.2]. The implication ($ii$) $\Rightarrow$ ($i$) is shown in Example 5.2. Finally, ($i$) $\Rightarrow$ ($iii$) follows from Theorem 5.6. $\square$
An interesting open question related to this result is whether the PD-approximation property for a Fell bundle ${\mathcal{B}}=(B_{g})_{g\in G}$ is equivalent to the AP, also in the case where $B_{e}$ is not nuclear.
6 Some further developments
In this section we briefly address some directions for further developments.
6.1
Let ${\mathcal{B}}=(B_{g})_{g\in G}$ be a Fell bundle over a discrete group $G$, and let $T$ be a ${\mathcal{B}}$-bundle map. It would be interesting to find some sufficient condition, other than $T$ being a linear combination of positive definite ${\mathcal{B}}$-bundle maps (cf. Remark 3.15), which ensures that $T$ is full or reduced, in which case one would also like to know whether $\Phi_{T}$ or $M_{T}$ is completely bounded. We mention here one result of this type. To any bounded function $\phi:G\to{\mathbb{C}}$ one may associate the ${\mathcal{B}}$-bundle map $T^{\phi}=(T^{\phi}_{g})_{g\in G}$ given by
$T^{\phi}_{g}(b)=\phi(g)b$
for every $g\in G$ and $b\in B_{g}$, as we did in Example 4.5 in the case where $\phi$ is positive definite. One may wonder under which conditions $T^{\phi}$ is full or reduced.
Let $B(G)$ denote the Fourier-Stieltjes algebra of $G$ [36], which consists of the coefficient functions of unitary representations of $G$. If $\phi$ belongs to $B(G)$, then $T^{\phi}$ is both full and reduced (because $\phi$ may be written as a linear combination of positive definite functions on $G$). One may also consider coefficient functions of more general representations of $G$ on Hilbert spaces, as was initiated by De Cannière and Haagerup [21] in the case of group bundles. (Note that this will bring something new only when $G$ is non-amenable.) The following result extends [9, Proposition 4.2] (which is itself an extension of [21, Theorem 2.2]) to the setting of Fell bundles.
Proposition 6.1.
Let $v$ be a uniformly bounded representation of $G$ into the invertible bounded linear operators on some Hilbert space ${\mathcal{H}}$, and set $K:=\sup\{\|v(g)\|,g\in G\}<\infty$. Let $\xi_{1},\xi_{2}\in{\mathcal{H}}$ and define $\phi:G\to{\mathbb{C}}$ by
$\phi(g)=\langle\xi_{1},v(g)\xi_{2}\rangle\quad\text{for all }g\in G.$
Then $T^{\phi}$ is reduced. In fact, the associated map $M_{T^{\phi}}:C_{r}^{*}({\mathcal{B}})\to C_{r}^{*}({\mathcal{B}})$ is completely bounded, with $\|M_{T^{\phi}}\|_{\rm cb}\leq K^{2}\,\|\xi_{1}\|\|\xi_{2}\|$.
Proof.
It is not difficult to check that there is an adjointable operator $W$ on the Hilbert $B_{e}$-module ${\mathcal{H}}\otimes\ell^{2}({\mathcal{B}})$ satisfying
$W(\xi\otimes j^{\mathcal{B}}_{h}(c))=v(h)\xi\otimes j^{\mathcal{B}}_{h}(c)\quad\text{ for all }\xi\in{\mathcal{H}},h\in G\text{ and~}c\in B_{h}.$
It is also clear that $W$ is invertible. Then we have
$W(I_{\mathcal{H}}\otimes\lambda^{\mathcal{B}}_{g}(b))W^{-1}=v(g)\otimes\lambda^{\mathcal{B}}_{g}(b)$
for every $g\in G$ and $b\in B_{g}$. Indeed, for every $\xi\in{\mathcal{H}},h\in G\text{ and~}c\in B_{h}$, we have
$W(I_{\mathcal{H}}\otimes\lambda^{\mathcal{B}}_{g}(b))(\xi\otimes j^{\mathcal{B}}_{h}(c))=W(\xi\otimes j^{\mathcal{B}}_{gh}(bc))=v(gh)\otimes j^{\mathcal{B}}_{gh}(bc)$
$=(v(g)\otimes\lambda^{\mathcal{B}}_{g}(b))(v(h)\otimes j^{\mathcal{B}}_{h}(c))=(v(g)\otimes\lambda^{\mathcal{B}}_{g}(b))W(\xi\otimes j^{\mathcal{B}}_{h}(c)).$
For $\xi\in{\mathcal{H}}$, let $\theta_{\xi}:\ell^{2}({\mathcal{B}})\to{\mathcal{H}}\otimes\ell^{2}({\mathcal{B}})$ be the adjointable operator given by $\theta_{\xi}f=\xi\otimes f$ for all $f\in\ell^{2}({\mathcal{B}})$, whose adjoint is determined by $\theta_{\xi}^{*}(\xi^{\prime}\otimes f)=\langle\xi,\xi^{\prime}\rangle\,f$ for $\xi^{\prime}\in{\mathcal{H}}$ and $f\in\ell^{2}({\mathcal{B}})$. We may now define $M_{\phi}:{\mathcal{L}}_{B_{e}}(\ell^{2}({\mathcal{B}}))\to{\mathcal{L}}_{B_{e}}(\ell^{2}({\mathcal{B}}))$ by
$M_{\phi}(x)=\theta_{\xi_{1}}^{*}W(I_{\mathcal{H}}\otimes x)W^{-1}\theta_{\xi_{2}}.$
Then $M_{\phi}$ is completely bounded with $\|M_{\phi}\|_{\rm cb}\leq K^{2}\,\|\xi_{1}\|\|\xi_{2}\|$. Moreover, we have
$M_{\phi}(\lambda^{\mathcal{B}}_{g}(b))f=\theta_{\xi_{1}}^{*}W(I_{\mathcal{H}}\otimes\lambda^{\mathcal{B}}_{g}(b))W^{-1}\theta_{\xi_{2}}f=\theta_{\xi_{1}}^{*}(v(g)\otimes\lambda^{\mathcal{B}}_{g}(b))\theta_{\xi_{2}}f$
$=\theta_{\xi_{1}}^{*}(v(g)\xi_{2}\otimes\lambda^{\mathcal{B}}_{g}(b)f)=\langle\xi_{1},v(g)\xi_{2}\rangle\,\lambda^{\mathcal{B}}_{g}(b)f=\phi(g)\,\lambda^{\mathcal{B}}_{g}(b)f$
for all $g\in G,b\in B_{g}$ and $f\in\ell^{2}({\mathcal{B}})$. Thus, $M_{\phi}$, restricted to $C_{r}^{*}({\mathcal{B}})$, is a completely bounded extension of $\phi_{T^{\phi}}$. This shows that $T^{\phi}$ is reduced, with $M_{T^{\phi}}=M_{\phi}$. $\square$
6.2
In this subsection, $G$ and $H$ are discrete groups, $\varphi\in{\rm Hom}(G,H)$, and $\phi:G\to{\mathbb{C}}$ is a bounded function. We will say that $\phi$ is a reduced $(G,\varphi,H)$-multiplier (or simply a $\varphi$-multiplier, when $G$ and $H$ are clear from the context) if there exists a bounded linear map $M_{\varphi,\phi}:C^{*}_{r}(G)\to C^{*}_{r}(H)$ such that
$M_{\varphi,\phi}(\lambda_{G}(g))=\phi(g)\lambda_{H}(\varphi(g))\ ,$
for every $g\in G$. One can also define full $(G,\varphi,H)$-multipliers in a similar way, in terms of the existence a suitable bounded linear map $\Phi_{\varphi,\phi}:C^{*}(G)\to C^{*}(H)$. Of course, when $G=H$ and $\varphi={\rm id}_{G}$, we recover the usual notions of reduced and full multipliers of $G$ (see e.g. [7] and references therein); the map $M_{{\rm id}_{G},\phi}$ is then denoted $M_{\phi}$, while $\Phi_{{\rm id}_{G},\phi}$ is denoted by $\Phi_{\phi}$. In this case, it is well-known (and follows from Theorems 3.12 and 3.10) that the function $\phi$ is positive definite if and only if $\phi$ is a reduced (resp. full) multiplier and $M_{\phi}$ (resp. $\Phi_{\phi}$) is completely positive.
Let’s rephrase these notions in terms of the group bundles ${\mathcal{A}}=({\mathbb{C}}\times\{g\})_{g\in G}$ and =(×${\mathcal{B}}=({\mathbb{C}}\times\{h\})_{h\in H}$. Let $T^{\phi}=(T^{\phi}_{g})_{g\in G}$ be the ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map given by
[…]
Then it follows readily that $\phi$ is a reduced (resp. full) $(G,\varphi,H)$-multiplier if and only if the bundle map $T^{\phi}$ is reduced (resp. full). Moreover, under the natural identifications $C_{r}^{*}({\mathcal{A}})\simeq C_{r}^{*}(G)$ and $C_{r}^{*}({\mathcal{B}})\simeq C_{r}^{*}(H)$ (resp. $C^{*}({\mathcal{A}})\simeq C^{*}(G)$ and $C^{*}({\mathcal{B}})\simeq C^{*}(H)$), the map $M_{T^{\phi}}$ corresponds to $M_{\phi}$ (resp. $\Phi_{T^{\phi}}$ corresponds to $\Phi_{\phi}$) whenever $\phi$ is a reduced (resp. full) $(G,\varphi,H)$-multiplier.
Note that $T^{\phi}=(T^{\phi}_{g})_{g\in G}$ is positive definite if and only if for all $g_{1},\ldots,g_{n}\in G$ and $z_{1},w_{1},\ldots,z_{n},w_{n}\in{\mathbb{C}}$, we have
$\sum_{i,j=1}^{n}(w_{i},\varphi(g_{i}))\,T^{\phi}_{g_{i}^{-1}g_{j}}\big{(}(z_{i},g_{i})^{*}(z_{j},g_{j})\big{)}\,(w_{j},\varphi(g_{j}))^{*}\,\geq 0\text{ in }{\mathbb{C}}\times\{e\}$
that is,
$\sum_{i,j=1}^{n}w_{i}\bar{w}_{j}\phi(g_{i}^{-1}g_{j})\bar{z}_{i}z_{j}\geq 0\,.$
This is clearly equivalent to $\phi$ being a positive definite function on $G$. Hence, Theorem 3.10 gives that $\phi$ is positive definite if and only $\phi$ is a full $(G,\varphi,H)$-multiplier such that $\Phi_{\varphi,\phi}:C^{*}(G)\to C^{*}(H)$ is completely positive. Moreover, if ${\rm ker}(\varphi)$ is amenable, then Theorem 3.12, gives that this is also equivalent to $\phi$ being a reduced $(G,\varphi,H)$-multiplier such that $M_{\varphi,\phi}:C_{r}^{*}(G)\to C_{r}^{*}(H)$ is completely positive.
The natural problem of determining the space of all reduced (resp. full) $(G,\varphi,H)$-multipliers, and eventually obtain some structure theorem, seems to be challenging. We include below a few relevant observations.
We first note that $\phi$ is a full multiplier of $G$ (i.e., $\phi$ belongs to $B(G)$) if and only if $\phi$ is a full $(G,\theta,K)$-multiplier for every group $K$ and every $\theta\in{\rm Hom}(G,K)$, in which case we have
$\Phi_{\theta,\phi}=\theta_{*}\circ\Phi_{\phi},$
where $\theta_{*}:C^{*}(G)\to C^{*}(K)$ is the canonical *-homomorphism, cf. [11, Corollary 8.C.13]. It would be interesting to know if there are situations where some full $(G,\varphi,H)$-multiplier is not a full multiplier of $G$.
Similarly, $\phi$ is a reduced multiplier of $G$ if and only if it is a reduced $(G,\theta,K)$-multiplier for every group $K$ and every $\theta\in{\rm Hom}(G,K)$ such that $\ker(\theta)$ is amenable, in which case we have
$M_{\theta,\phi}=\theta_{*,r}\circ M_{\phi},$ (7)
where $\theta_{*,r}:C_{r}^{*}(G)\to C_{r}^{*}(K)$ is the canonical *-homomorphism [11, Corollary 8.C.15].
If $\ker(\varphi)$ is not amenable, then it can happen that $\phi$ is a reduced multiplier of $G$, but not a reduced $(G,\varphi,H)$-multiplier. For example, assume $G$ is a non-amenable group, $H=\{e\}$ and $\varphi_{0}:G\to\{e\}$ is the trivial map. Then $\ker(\varphi_{0})$ is not amenable. Moreover, the constant function $1_{G}$ is a reduced multiplier of $G$, but it is not a reduced $(G,\varphi_{0},\{e\})$-multiplier. Indeed, if it was, then $M_{\varphi_{0},1_{G}}$ would correspond to a multiplicative linear functional on $C_{r}^{*}(G)$ taking the value $1$ on each $\lambda_{G}(g)$, which would imply that $G$ is amenable, cf. [14]. It should also be noted here that the space of reduced $(G,\varphi_{0},\{e\})$-multipliers may be identified with the dual of $C_{r}^{*}(G)$.
One may wonder whether there exists some reduced $(G,\varphi,H)$-multiplier which is not a reduced multiplier of $G$, even if $\ker(\varphi)$ is amenable.
Let now $\psi:H\to{\mathbb{C}}$ be a bounded function, so $\psi\circ\varphi:G\to{\mathbb{C}}$ is bounded too. We note that if $\psi$ is a full multiplier of $H$, i.e., $\psi\in B(H)$, then $\psi\circ\varphi$ belongs to $B(G)$, cf. [36, Theorem 2.2.1 (i)]; hence $\psi\circ\varphi$ is a full $(G,\varphi,H)$-multiplier, satisfying that
$\Phi_{\varphi,\psi\circ\varphi}=\Phi_{\psi}\circ\varphi_{*}=\varphi_{*}\circ\Phi_{\psi\circ\varphi}.$
We don’t know if it can happen that $\psi\circ\varphi$ is a full $(G,\varphi,H)$-multiplier for some $\psi\not\in B(H)$.
Further, assume that $\ker(\varphi)$ is amenable and let $\psi$ be a reduced multiplier of $H$. Then one readily sees that $\psi\circ\varphi$ is a reduced $(G,\varphi,H)$-multiplier with
$M_{\varphi,\psi\circ\varphi}=M_{\psi}\circ\varphi_{*,r}.$ (8)
It is not obvious to us that $\psi\circ\varphi$ is necessarily a reduced multiplier of $G$. However, if $M_{\psi}$ is completely bounded, i.e., $\psi\in M_{0}A(H)$, then $\psi\circ\varphi\in M_{0}A(G)$ (this may be deduced from [35, Theorem], which is a variant of Gilbert’s theorem), in which case we get that
$M_{\varphi,\psi\circ\varphi}=M_{\psi}\circ\varphi_{*,r}=\varphi_{*,r}\circ M_{\psi\circ\varphi}.$
Also here, one may wonder whether it can happen that $\psi\circ\varphi$ is a reduced $(G,\varphi,H)$-multiplier for some $\psi$ which is not a reduced multiplier of $H$.
We finally note that if $\psi$ is assumed to be positive definite, then it follows readily that the maps $\Phi_{\varphi,\psi\circ\varphi}$ and $M_{\varphi,\psi\circ\varphi}$ are both completely positive.
6.3
There has been recently a lot of interest in $C^{*}$-algebras associated to Fell bundles, not only over discrete groups but also over locally compact groups, étale groupoids and inverse semigroups (see e.g. [15, 1, 17, 41, 2, 42, 43, 39, 44, 38, 20, 61]). It seems natural to investigate to which extent our results in this paper can be carried over to these settings. Although some work will be necessary in order to handle some technicalities, we expect that most of our results will continue to hold for Fell bundles over locally compact groups. In the other cases, it is likely that some efforts will be needed to formulate the correct versions. For instance, for étale groupoids, one would first have to decide which class(es) of functors between étale groupoids have the property that a suitable generalization of [11, Theorems 8.C.12 and 8.C.14] is achievable. We refer to [61] for some possible candidates.
In another direction, it has been known for quite some time that Fell bundles over (discrete) groupoids and $C^{*}$-categories are close relatives, although it is hard to find any specific statement in the literature. Let ${\cal T}$ be a (small) $C^{*}$-category [30]. One can naturally associate to ${\cal T}$ a Fell bundle ${\mathcal{B}}_{\cal T}$ over the (discrete) pair groupoid ${\rm Ob}({\cal T})\times{\rm Ob}({\cal T})$ by setting $B_{\rho,\sigma}:=(\rho,\sigma)$, the Banach space of arrows from the object $\rho$ to the object $\sigma$ in ${\cal T}$. The bundle operations follow at once from the corresponding operations in ${\cal T}$. If $t\in B_{\rho,\sigma}$ and $s\in B_{\sigma,\tau}$ then $ts\in B_{\rho,\tau}$ (product in the bundle) is given by $s\circ t$ (composition in ${\cal T}$), and $t^{*}\in B_{\sigma,\rho}$ (adjoint in the bundle) is nothing but $t^{*}\in(\sigma,\rho)$. Moreover, it is well known that to any $C^{*}$-category ${\cal T}$ one can associate a $C^{*}$-algebra $C^{*}({\cal T})$ (cf. [30]), and to any Fell bundle ${\mathcal{B}}$ over a groupoid one can associate $C^{*}({\mathcal{B}})$ and $C_{r}^{*}({\mathcal{B}})$ (see e.g. [37]). One may wonder whether there is some relationship between $C^{*}({\cal T})$ and $C^{*}({\mathcal{B}}_{\cal T})$ (or $C_{r}^{*}({\mathcal{B}}_{\cal T})$).
Note also that if ${\cal T}$ is a monoidal (i.e., tensor) $C^{*}$-category, then a second operation of composition in the Fell bundle ${\mathcal{B}}_{\cal T}$ arises from the tensor product in ${\cal T}$, namely if $t\in B_{\rho,\sigma}$ and $t^{\prime}\in B_{\rho^{\prime},\sigma^{\prime}}$ then we get $t\otimes t^{\prime}\in B_{\rho\otimes\rho^{\prime},\sigma\otimes\sigma^{\prime}}$ (we leave the reader the task of spelling out the detailed properties of this composition). Thinking of a monoidal $C^{*}$-category as a special case of a 2-$C^{*}$-category with only one object (see e.g. [62]), this might open the way to a parallel notion of $2$-Fell bundle (cf. [12]).
Over the years, notions of amenability have appeared in the context of $C^{*}$-categories mainly inspired by subfactor theory, see e.g. [48] for the concept of amenability of an object in a monoidal $C^{*}$-category. In the more recent paper [57], many properties originally defined for groups (e.g., amenability, property (T), the Haagerup property and weak amenability) are translated into the setting of rigid monoidal $C^{*}$-categories. It is therefore natural to ask if the amenability of a rigid monoidal $C^{*}$-category ${\cal T}$ is related to some suitable notion of amenability (or approximation property) of the corresponding Fell bundle ${\mathcal{B}}_{\cal T}$. It should be possible to put forward several properties from groups to Fell bundles, and we believe that the material developed in the main text should provide valuable tools for this purpose.